### v2

# David Roodman Open Philanthropy San Francisco, CA

Abstract
Like Python and Java, which are integrated into Stata, Julia is a free programming language that runs on all major operating systems. The julia package links Stata to Julia as well. Users can transfer data between Stata and Julia at high speed, issue Julia commands from Stata to analyze and plot, and pass results back to Stata. Julia’s econometric software ecosystem is not as mature as Stata’s or R’s, or even Python’s. But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) can call a Julia back end for a 33–50% speed-up, even as the R package fwildclusterboot (Fischer and Roodman, 2021) uses the same back end for inference after instrumental variables estimation. reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls an independently developed Julia package for tenfold acceleration on hard problems. reghdfejl also supports nonlinear models—preliminarily, as the Julia package for that purpose matures.
NOTE: This article is formatted for submission to the Stata Journal. It has not been accepted.
keywords:
st0001, reghdfe, reghdfejl, boottest, Julia, high-dimensional fixed effects, cross-platform communication
1 Introduction
Whenever you issue a command to perform a statistical calculation, you initiate a chain of events of bewildering complexity: your request cascades through layers of code crafted over decades by thousands of programmers in multiple languages, as it is expanded into millions or billions of computational steps—steps that are executed, if you’re lucky, in the time it takes to read this sentence.
In Stata, for example, a call to regress or gsem starts out in the ado programming language. The ado code translates your command line into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language. In modern Stata, an ado program can even call Python or Java. The called code will probably invoke the LAPACK library, which is written in Fortran, to invert matrices or solve linear systems.
Here I introduce a package that forges a fresh link between Stata and another language, Julia. Julia is new: it reached version 1.0 in 2018. It is also free, and available for Windows, Linux, and macOS. Like Python code, Julia code is fully portable across those platforms and can run as soon as downloaded.
For creators of computationally intensive software, Julia has many strengths. More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it. Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution. (Python and C++ is a common combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that code for reuse. Like LISP, a language almost as old as Fortran, Julia can treat code as data. This feature opens the door to meta-programs that can, for example, automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
Julia has disadvantages. While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented. Standards for passing results are less well developed than in Stata. Much less statistical functionality is available in Julia than in Stata or R, or even Python.
Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development. Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms. The core work can implemented, optimized, and augmented in one place, rather than separately on each platform.
The julia package links Stata to Julia. It follows in the footsteps of many projects to link statistical software to other languages. StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003. It added the built-in Mata language in Stata 9. Fiedler (2012) envisioned running Python inside of Stata. Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a plugin for Stata. Perhaps inspired by Fiedler, Stata 16 fully integrated Python. Java plugins became possible in Stata 13 and the Java prompt debuted in the Stata results window in version 17. Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk. Outside of Stata, the JuliaCall package (Li, 2019) lets R users call Julia routines. An identically named package does the same for Python, and is paired with PyCall for Julia. [^1] See juliapy.github.io/PythonCall.jl/stable.
The julia package was most inspired by Fiedler (2013). Its core too is a C++ plugin, which comes precompiled for Windows, Linux, macOS/Intel, and macOS/ARM. From the user’s point of view, the package has three components:
  • The jl: command for Stata, which prefixes single Julia lines or, when typed by itself (with or without the colon), starts a Julia session in Stata.
  • jl subcommands for managing Julia packages and transferring data between Stata and Julia.
  • Julia functions to read and write Stata macros, scalars, matrices, and variables.
The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package.
Two user-written Stata packages call Julia via julia. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata back end. And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016), while calling FixedEffectModels.jl, whose lead developer is Matthieu Gomez. reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl. In the last case, using Julia brings new functionality to Stata. In all cases, calling Julia cuts runtime on hard problems, sometimes by an order of magnitude.
Stata users can thus benefit from Julia and julia without using them directly. Still, this paper explains the new julia package. So it starts with a hands-on demonstration of running Julia from Stata, in section 2. Then section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia. Section 4 gives tips and warns of some limitations. Section 5 more thoroughly describes the package’s Stata commands and Julia functions.
2 Hands-on introduction
2.1 Installation and first examples
Julia is an interpreted language in the sense that you can only run programs written in it if you have installed it. [^2] PackageCompiler.jl can build stand-alone Julia apps, but the files take hundreds of megabytes because they contain complete copies of Julia and its libraries. In contrast, compiled C++ and Fortran programs run on computers without a C++ or Fortran compiler.
The preferred way to get Julia is to download the juliaup installation manager. You do that in Windows by installing “Julia” (not “juliaup”) from the Windows Store. [^3] The instructions might change. See github.com/JuliaLang/juliaup#installation for the latest. In Linux and macOS, you can instead start Stata and type
  • ! curl -fsSL https://install.julialang.org | sh -s -- -y
(including the “!” at the beginning). In Linux, you should then restart Stata and the shell to trigger the addition of Julia’s location to the PATH environment variable.
To install the julia package in Stata, type:
  • ssc install julia
You can now call Julia from Stata:
  • . jl: "Hello world!"
    Starting Julia
    "Hello world!"
The output may take several minutes to appear. On first use, jl installs and precompiles two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on. DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets—i.e., rectangular grids whose columns have names and potentially different data types. CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
In general when using Julia, you will experience delays at four points: when packages are installed; when they are first used; when they are first used in a Julia session; and when particular features within them are first used, such as clustering standard errors. A thrust of recent development in the Julia langauge has been to move the behind-the-scenes processing from later to earlier stages in order to shorten the lags during regular use.
In the session shown below, I load Stata’s venerable “auto” data set, extract five variables, run a regression on them, and then copy the data to a Julia DataFrame named auto:
  • . sysuse auto
    (1978 automobile data)
    . keep price headroom foreign mpg turn
    . reg price headroom foreign#c.mpg, cluster(turn)
    Linear regression                               Number of obs     =         74
                                                    F(3, 17)          =       9.69
                                                    Prob > F          =     0.0019
                                                    R-squared         =     0.2939
                                                    Root MSE          =       2531
                                       (Std. err. adjusted for 18 clusters in turn)
                                  Robust
            price   Coefficient  std. err.      t    P>|t|     [95% conf. interval]
         headroom    -273.2403   274.4352    -1.00   0.333     -852.248    305.7674
    foreign#c.mpg
        Domestic     -344.0717   72.40801    -4.75   0.000    -496.8392   -191.3041
         Foreign     -267.0448   57.09596    -4.68   0.000    -387.5067   -146.5828
            _cons     13743.64     2176.8     6.31   0.000      9150.99    18336.28
    . jl save auto
    Data saved to DataFrame auto in Julia
jl save is not a Julia command. It is a subcommand of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten.
Next I type jl by itself, which changes the prompt in much the same way that typing mata or python or java does. Have started an interactive Julia session, I display the new DataFrame by typing its name:
  • . jl
                             Julia (type exit() to exit)
    jl> . auto
    74×5 DataFrame
     Row   price   mpg     headroom  turn    foreign
           Int16?  Int16?  Float32?  Int16?  Cat…?
       1     4099      22       2.5      40  Domestic
       2     4749      17       3.0      40  Domestic
       3     3799      22       3.0      35  Domestic
       4     4816      20       4.5      40  Domestic
       5     7827      15       4.0      43  Domestic
       6     5788      18       4.0      43  Domestic
       7     4453      26       3.0      34  Domestic
       8     5189      20       2.0      42  Domestic
      68     3748      31       3.0      35  Foreign
      69     5719      18       2.0      36  Foreign
      70     7140      23       2.5      36  Foreign
      71     5397      41       3.0      35  Foreign
      72     4697      25       3.0      35  Foreign
      73     6850      25       2.0      36  Foreign
      74    11995      17       2.5      37  Foreign
                                           59 rows omitted
Along the top are the column names, and below them the types. Int16 means a 16-bit integer, equivalent to Stata int. Float32 corresponds to Stata float. The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing. By default, jl save made all the columns accept missing. jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector. While the Julia version of the variable also has an internal, integer representation, the vector presents as a text variable. [^4] Julia categorical variables are not quite comparable to Stata variables with data labels. Code involving the Julia variable refers to display values, such as “Domestic,” not to numerical codes. The type label “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
Next I run the same regression in Julia, with the reg() function from the Julia package FixedEffectModels.jl:
  • jl> . using FixedEffectModels
    jl> . reg(auto, term(:price) ~ term(:headroom) + term(:foreign) & term(:mpg),
      ...     cov.cluster(:turn))
    FixedEffectModel
    Number of obs:                           74   Converged:                             true
    dof (model):                              3   dof (residuals):                         16
    R²:                                   0.294   R² adjusted:                          0.264
    F-statistic:                         9.6942   P-value:                              0.001
                              Estimate  Std. Error     t-stat  Pr(>|t|)  Lower 95%  Upper 95%
    headroom                  -273.24      274.435  -0.995646    0.3342   -855.017    308.536
    foreign: Domestic & mpg   -344.072      72.408  -4.75185     0.0002   -497.57    -190.574
    foreign: Foreign & mpg    -267.045      57.096  -4.67712     0.0003   -388.083   -146.007
    (Intercept)              13743.6      2176.8     6.31369     <1e-04   9129.03   18358.2
The Stata and Julia results match. But that Julia regression command needs explaining. First some cosmetic comments:
  • Because the regression command was long, I broke it in two. After I hit the enter key at the end of the first line, jl detected that the line was not grammatical by itself and printed “...” to prompt me to continue.
  • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears here.
  • All variable names in the command line are prefixed with “:”. That indicates that the names are symbols, not stand-alone Julia objects like the DataFrame auto.
  • Because foreign is a CategoricalVector, reg() automatically treats it as a factor variable, generating a pair of dummies to be interacted with mpg. There is no need for an “i.” prefix or the equivalent as in Stata. However, treating non-CategoricalVectors as factor variables requires a more awkward syntax. One must pass reg() an argument with a dictionary mapping the variables to coding types, such as with contrasts=Dict(:headroom=>DummyCoding()).
The command line contains deeper complexity. In Julia, as in R and Python, every regression command is a statement in a full-fledged programming language. This tends to complicate the syntax for even basic operations. It is hard to implement an R or Python line as clean as regress y x. In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model. The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and surfaces for the user. The content of each Term is just a symbol, such as :price. What give Term’s life are the operations that StatsModels.jl defines on them. The + operator is repurposed (“overloaded”) to form tuples of Term’s—pairs, triplets, and so on. The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides. The entire FormulaTerm is passed as the second argument to reg(). In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
This standard for representing regression models is not intrinsic to Julia. It has emerged from the work of users. And it is transparent. In Stata, one could type jl: dump(term(:c) $\sim$ term(:a) + term(:b)) to reveal the contents of a FormulaTerm. The software that builds and applies these structures is on GitHub, all written in Julia. The StatsModels.jl formula language is also extensible [^5] See juliastats.org/StatsModels.jl/stable/internals/#extending.. As we will see, FixedEffectModels.jl defines a way to mark variables whose fixed effects are to be absorbed.
Yet a powerful feature of Julia makes it practical to ignore most of this complexity. The regression command above can be simplified with a macro. Where macros in Stata are strings, in Julia they are code for manipulating code—meta-programs. StatsModels.jl defines @formula, which lets one write models in a more R-like way. Thus:
  • jl> . m=reg(auto, @formula(price ~ headroom + foreign & mpg), Vcov.cluster(:turn))
    [identical output omitted]
Before this new command is compiled and executed, @formula turns it into something closer to the previous one. [^6] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones. An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
Notice one other change to the command line: it now starts with m= in order to store the fitted model in the Julia variable m. This lets me extract return values and pass them back to Stata:
  • jl> . st_numscalar("adjR2", adjr2(m))
    jl> . exit()
    . display "Adjusted R2 = " adjR2
    Adjusted R2 = .26365506
The Julia function st_numscalar() is part of the julia package; like its Mata namesake, it reads and writes Stata scalars.
2.2 Built for speed
To partially convey why Julia is a good environment for developing numerical software, I present another example. Suppose we have a data set with 10 million rows and 10 columns named x1-x10. Given a $10\times 10$ matrix ${\bf Q}$, we want to compute the norm of each row according to the quadratic form defined by ${\bf Q}$ and store the results in a new column. That is, for each row $x_{i}$, now viewed as a column vector, we want $x_{i}^{\prime}{\bf Q}x_{i}$. It is natural to implement this computation with triply nested for loops. That’s a good way to do it in Julia (inside Stata):
  • jl AddPkg LoopVectorization  // one-time package installation from GitHub
    jl: using LoopVectorization  // load package for use now
    jl: function XQX(Q, X)                             ///
          N, M = size(X);                              ///
          retval = zeros(N,1);                         ///
          @tturbo for i in 1:N                         ///
            for j in 1:M                               ///
              for k in 1:M                             ///
                retval[i] += X[i,j] * Q[j,k] * X[i,k]  ///
              end                                      ///
            end                                        ///
          end;                                         ///
          return retval                                ///
        end
Unlike in Python, indentation has no semantic meaning; this code snippet is indented only for readability.
We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()). But the interactive mode is not accessible to do files. To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line. It is typographically split across multiple lines with Stata’s continuation token, ///. When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon.
The code for this new Julia function, XQX(), is mostly straightforward. It initializes the return value to a column of 0’s, performs the calculation through nested loops, and returns the result. One detail is unusual: another macro is invoked. @tturbo is part of Chris Elrod’s LoopVectorization.jl package. @tturbo analyzes the code that follows it and substantially—but invisibly—rewrites it for speed. It may unroll a loop to reduce the number of jumps back to the top. The macro, or the compiler, may spot the opportunity to move X[i,j] * out of the inner loop since it does not depend on the loop’s index, k. The macro may even reorder the nesting of the loops with an eye toward the fact that in Julia, matrices are stored column by column, and on-chip memory caching makes it is faster to access adjacent memory locations in sequence. The LoopVectorization package is called that because it also tailors code so that the resulting machine language exploits the ability of modern chips to vectorize operations within each core, in 256- or 512-bit registers—what is called single instruction, multiple data (SIMD) execution. And @tturbo will exploit multithreading, the subdivision of work across cores; that is what the first “t” in tturbo stands for.
Triply nested implementations in Mata and Python are much slower, I assume because the calculations in the innermost loop have to be interpreted a billion times. The best alternatives I have found in the two languages are essentially the same:
  • mata
    function XQX(Q, X)
      return (rowsum((X * Q) :* X))
    end
    python
    def XQX(Q,X):
      return ((X @ Q) * X).sum(axis=1)
    end
Both delegate the computation to built-in matrix operations, which are fast because they are implemented in compiled C or Fortran. However, both contain a subtle inefficiency. They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$. Since in our example ${\bf X}$ is $10^{6}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each. Allocating and deallocating memory itself takes time. Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory. The triply-looped Julia code demands essentially no temporary storage.
To create a test bed for these functions, I construct ${\bf X}$ as a Stata data set and ${\bf Q}$ as a Stata matrix, then copy them into Mata, Python, and Julia. The set rmsg on command generates timings:
  • . python: import numpy as np
    . python: from sfi import Data, Matrix
    . set obs 10000000
    Number of observations (_N) was 0, now 10,000,000.
    . drawnorm double x1-x10
    . mata  : st_matrix("Q", makesymmetric(runiform(10,10)))
    . mata  : Q = st_matrix("Q")
    . python: Q = np.asarray(Matrix.get("Q"))
    . jl    : Q = st_matrix("Q");
    . set rmsg on
    . mata: X = st_data(.,"x1-x10")
    r; t=0.07
    . python: X = np.asarray(Data.get(var="x1 x2 x3 x4 x5 x6 x7 x8 x9 x10"))
    r; t=6.73
    . jl: X = st_data("x" .* string.(1:10));
    r; t=0.24
    . jl: X = st_data("x" .* string.(1:10));
    r; t=0.15
Under a Julia convention, the semicolons at the end of the jl lines suppress the printing of the results of the assignments. The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because of just-in-time compilation.
On a Windows laptop with an Intel i9-13900H, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python. In jl, the data copying is performed by C++ code that is multithreaded across columns.
Next I test the three functions I defined:
  • . mata: y = XQX(Q,X)
    r; t=0.31
    . python: y = XQX(Q,X)
    r; t=0.48
    . jl: y = XQX(Q,X);
    r; t=1.39
    . jl: y = XQX(Q,X);
    r; t=0.05
The Julia function is especially slow on first use because that is when @tturbo performs its magic. On later calls, Julia is fast: it takes 0.05 seconds, which works out to about 40 billion floating point operations per second. Mata takes 0.31 seconds and Python, 0.48. To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above. In order to trigger immediate compilation, the Julia package manager would run the script after installing the package. Julia would save the generated machine code for future use.
3 Applications
This section presents three applications of julia. They involve HDFE modeling, wild bootstrap–based inference, and plotting.
3.1 Estimation with high-dimensional fixed effects
The introduction of reghdfe in 2014 (Correia, 2016) was an important event in applied econometrics. Building on Guimarães and Portugal (2010) and Gaure (2011), the Mata-based program proved the computational tractability of linear models with many sets of fixed effects. Since then, other authors have developed alternative algorithms for absorbing fixed effects, several of which have been incorporated into reghdfe. The ideas have been extended to the Poisson model through ppmlhdfe for Stata (Correia, Guimarães, and Zylkin, 2019), as well as to other generalized linear models, in alpaca and fixest for R (Stammann 2018; Bergé 2018). Julia packages for HDFE modeling are also available: Matthieu Gomez’s FixedEffectModels.jl and Johannes Boehm’s GLFixedEffectModels.jl. With regard to linear models, according to benchmarking results on the GitHub pages of fixest and FixedEffectModels.jl, those two packages are 1–2 orders of magnitude faster than regdhfe. [^7] See github.com/lrberge/fixest#benchmarking and github.com/FixedEffects/FixedEffectModels.jl#benchmarks.
These developments have given R and Julia users faster and more general implementations of HDFE modeling than were available in Stata. That is what motivated the development of the julia and reghdfejl packages. The idea was to not reinvent wheels, but to fashion a familiar-looking chassis onto existing wheels.
The extended example in section 2.1 followed a sequence that is the essence of reghdfejl: copy data from Stata to Julia, estimate in Julia, return results to Stata. Here I complete the narrative arc of that example. With a small typographic change to the last Julia estimation line—wrapping a variable name in fe()—I exploit the ability of FixedEffectModels.jl to absorb fixed effects:
  • . jl: reg(auto, @formula(price ~ headroom + fe(foreign) & mpg), Vcov.cluster(:turn))
    FixedEffectModel
    Number of obs:                     74  Converged:                       true
    dof (model):                        1  dof (residuals):                   34
    R²:                             0.294  R² adjusted:                    0.264
    F-statistic:                 0.991311  P-value:                        0.334
    R² within:                      0.583  Iterations:                         1
                 Estimate  Std. Error     t-stat  Pr(>|t|)  Lower 95%  Upper 95%
    headroom      -273.24     274.435  -0.995646    0.3342   -855.017    308.536
    (Intercept)  13743.6     2176.8     6.31369     <1e-04   9129.03   18358.2
Back in Stata, here is reghdfejl running the same regression:
  • . reghdfejl price headroom, absorb(foreign#c.mpg) cluster(turn)
    (MWFE estimator converged in 1 iterations)
    HDFE linear regression with Julia                 Number of obs   =         74
    Absorbing 1 HDFE group                            F(   2,     15) =       0.99
    Statistics cluster-robust                         Prob > F        =     0.3941
                                                      R-squared       =     0.2939
                                                      Adj R-squared   =     0.2637
    Number of clusters (turn)    =         18         Within R-sq.    =     0.5829
                                                      Root MSE        =  2530.9785
                                     (Std. err. adjusted for 18 clusters in turn)
                 |               Robust
           price | Coefficient  std. err.      t    P>|t|     [95% conf. interval]
        headroom |  -273.2403   274.4352    -1.00   0.334     -855.017    308.5364
           _cons |   13743.64     2176.8     6.31   0.000     9129.026    18358.24
In spirit, reghdfejl is a slot-in replacement for reghdfe, with the same syntax, output, and e() return values. In practice, the two differ outwardly as well as inwardly. As reghdfjle depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience. reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering. So it often reports slightly larger standard errors. While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007). As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic. reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and pass the results to ivreg2. But in this mode, reghdfejl is no faster than ivreghdfe. For on hard problems, ivreg2 dominates the runtime.
reghdfe and reghdfejl share the following options: absorb(), vce(), residuals(), tolerance(), iterate(#), nosample, keepsingletons, compact, level(#), and standard Stata regression display options. In particular, reghdfejl’s absorb() can save some or all of the estimated fixed effects. For instance, a(firm year_fe=year) will absorb firm and year fixed effects, and store estimates of the latter in the variable year_fe. For both reghdfe and reghdfejl, the tolerance() option sets a precision threshold for the fitting algorithms to declare convergence; however, the semantics are not necessarily the same, so that, say, tol(1e-6) could elicit slightly different results from the two.
reghdfejl accepts three distinctive (documented) options:
gpu requests the use of a GPU. Currently the option saves more time on NVIDIA than Apple GPUs.
threads(#) lowers the limit on the number of threads that can be used. It cannot increase the thread count above the limit set when Julia is started (see section 4.2). This option is rarely used.
verbose directs reghdfejl to show more of its work—to display the Julia copy of the data set, the formula for the regression model, and the regression command. The data set and formula are left behind for the user to work with through jl, not erased as they otherwise would be.
In addition, reghdfejl offers a bootstrap variance-covariance estimator. vce(bootstrap) (or vce(bs)) is standard within Stata; regress accepts it. As an example, for reghdfejl, the clause
  • vce(bs, reps(1000) seed(42) procs(4) cluster(year))
specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from clusters defined by year. One could attain the same results with bootstrap:reghdfe or bs:reghdfe. But reghdfejl implements vce(bs) in a radically faster way. After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times, etc.).
To compare reghdfe and reghdfejl on speed, I create a data set with 10 million rows and two sets of fixed effects. I run the same regression with both commands. I explore the effect of switching from one to multiple processors in Stata/MP, and of adding reghdfejl’s gpu option:
  • set obs 10000000
    gen id1 = runiformint(1, 100000)
    gen id2 = runiformint(1, 100)
    drawnorm x1 x2
    gen double y = 3 * x1 + 2 * x2 + sin(id1) + cos(id2) + runiform()
    set processors 1
    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
    set processors 6
    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
For comparability, the reghdfe lines include dof(none) to disable the degrees-of-freedom correction, since reghdfejl lacks that feature. The first triplet runs on one CPU core, as under a non-MP flavor of Stata. But that setting on restricts the ado and Mata code: the Julia code still runs on multiple CPU threads. On a laptop with an Intel i9-13900H CPU and an NVIDIA RTX 4070 GPU, reghdfe takes 32 seconds. reghdfejl takes 3.8 seconds without the gpu option and 2.8 with. Moving to 6 processors in Stata and Mata—since the Intel chip has 6 “performance” cores—lowers the times to 24, 3.3, and 2.7 seconds.
Another script benchmarks reghdfejl’s vce(bs) option:
  • webuse nlswork
    xtset, clear
    set processors 6
    bs, cluster(occ_code) reps(1000): ///
      reghdfe ln_wage age ttl_exp tenure not_smsa south, absorb(year occ_code) dof(none)
    parallel initialize 6
    parallel bs, cluster(occ_code) reps(1000): ///
      reghdfe ln_wage age ttl_exp tenure not_smsa south, absorb(year occ_code) dof(none)
    reghdfejl ln_wage age ttl_exp tenure not_smsa south, absorb(year occ_code) ///
      vce(bs, cluster(occ_code) reps(1000) procs(6))
Using Stata’s bs prefix command with reghdfe takes 64 seconds. Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 6 copies of Stata [^8] See github.com/gvegayon/parallel. cuts the time to 23 seconds. reghdfejl takes just 2.5 seconds. But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 6 worker copies of Julia and triggers some just-in-time compilation in each. The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers is kept at 6.
The reghdfejl package includes a partialhdfejl command for absorbing fixed effects out of other variables. partialhdfejl command lines must contain either a generate() option to name the variables that will hold the results, or a prefix() option for making the new variable names out of the old. If any of the new variables already exists, partialhdfejl will error, unless the replace option is also included. The following two runs produce the same point estimates:
  • reghdfejl ln_wage age ttl_exp tenure south, absorb(year occ_code) cluster(occ_code)
    partialhdfejl ln_wage age ttl_exp tenure south, absorb(year occ_code) prefix(_)
    regress _ln_wage _age _ttl_exp _tenure _south, cluster(occ_code) nocons
To fit nonlinear HDFE models with reghdfejl, one includes family() and/or link() options in the command line, following the syntax of the built-in Stata command glm. This code below fits an HDFE Poisson model with the Mata-based ppmlhdfe as well as with reghdfejl. The latter runs about twice as fast:
  • use http://fmwww.bc.edu/RePEc/bocode/e/EXAMPLE_TRADE_FTA_DATA
    egen imp = group(isoimp)
    egen exp = group(isoexp)
    expand 100  // expand data set 100-fold for tougher test
    ppmlhdfe  trade fta, a(imp#year exp#year imp#exp) cluster(imp#exp)
    reghdfejl trade fta, a(imp#year exp#year imp#exp) cluster(imp#exp) family(poisson)
In the same way, one can fit HDFE logit, binomial, negative binomial, and other models. However, at this writing, the underlying Julia package, GLFixedEffects.jl, lacks key features such as observation weighting and the equivalent of Stata’s exposure() option for Poisson models. So reghdfejl’s ability to fit nonlinear models is currently undocumented.
3.2 Wild bootstrapping
Roodman et al. (2019) introduces the Stata package boottest for inference via the wild bootstrap. Among the methods boottest implements is the “wild restricted efficient” bootstrap (WRE), which is the Davidson and MacKinnon (2010) adaptation for instrumental variables (IV) estimation. Since the publication of Roodman et al. (2019), boottest has been optimized in ways that especially benefit the WRE. Using the Frisch-Waugh-Lowell theorem, exogenous controls are now partialled out of the other variables using the same algebraic tricks that accelerate boottest’s wild bootstrap for ordinary least squares. The partialling-out accelerates the application of the user’s IV estimator despite its irreducible nonlinearity (Roodman et al., 2019, section 6.1). Also incorporated (more fully), is MacKinnon (2023)’s point that the data only enter wild-bootstrapped inferences through indexed sets of matrix products such as $\{{\bf X}^{\prime}_{g}{\bf Y}_{g}:g=1,...,G\}$ where ${\bf X}_{g}$ and ${\bf Y}_{g}$ hold the rows of data matrices ${\bf X}$ and ${\bf Y}$ for cluster $g$. Reducing the data to these products early in the calculation complicates the algebra the code must implement, but speeds execution when clusters are few.
Finally, the back-end functionality in boottest has been transplanted to the Julia package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end.
Section 8.4 of Roodman et al. (2019) approximately replicates two 2SLS regressions in the Levitt (1996) study of the short-term impact of decarceration on crime. [^9] Data and preparation code are available at davidroodman.com/david/Levitt.zip. Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for adding the julia option to the boottest command line:
  • use Levitt, clear
    set seed 8723419
    foreach crimevar in Violent Property {
      reghdfejl D.l`crimevar'pop (LD.lpris_totpop = ibnL.stage#i(1/3)L.substage) ///
        D.(lincomepop unemp lpolicepop metrop black a*pop) i.year, a(state) clust(state year)
      boottest LD.lpris_totpop,bootclust(year) ptype(equaltail) ///
        gridmin(-2) gridmax(2) nograph
On a single CPU core, version 2.3.5 of boottest, the one published with Roodman et al. (2019), needs 210 seconds to perform the two boottest calls. The latest version finishes in 7.4 seconds. Adding the julia option to the boottest line brings the time to 4.1 seconds.
The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE. This may be the first instance of a Julia statistical package serving as a back end on multiple platforms.
3.3 Plotting
Many data visualization packages are available for Julia. Surveying them is beyond the scope of this article. To indicate their potential for Stata users, I generate two graphs with Makie.jl.
Makie works with several plotting back ends, which render output in formats that are optimized for different applications, such as interactive graphs in web pages and publication-quality image files. The Cairo back end is best for the latter. It is bundled with Makie in CairoMakie.jl:
  • jl AddPkg CairoMakie  // one-time installation
    jl: using CairoMakie   // use now
In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.101010blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command. margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights. The probabilities are copied to Python and depicted in a three-dimensional surface plot. A Julia adaptation of that example is:
  • webuse nhanes2
    svy: logistic highbp c.age##c.weight
    margins, at(age=(20(5)80) weight=(40(5)180))
    matrix xyz = r(at), r(b)'  // x, y, z values in 3 columns
    jl
      df = DataFrame(st_matrix("xyz"), [:age, :weight, :pr_highbp])  # get data
      f = surface(df.age, df.weight, df.pr_highbp,
                  axis=(type=Axis3,
                        title = "Probability of Hypertension by Age and Weight",
                        xlabel = "Age (years)",
                        ylabel = "Weight (kg)",
                        zlabel = "Probability of Hypertension"))
      f |> display  # equivalent to display(f)
      save("fig1.png", f)
    exit()
This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines. The first Julia line imports the data as a matrix and then converts it to a DataFrame with named columns. Then the surface plot is built and stored in the Julia variable f. The call to surface() refers to columns of the DataFrame with the convenient dot syntax, as in df.age; here, column names are not prefixed with a colon. The next command displays the figure in a new window, which might pop up behind the Stata window. That line takes advantage of Julia’s piping syntax for function calls. Last, the image is saved to disk. It appears in Figure 1.
Figure 1: Contour plot made with Makie in Julia
From the same data set, I plot the distribution of body mass index (BMI) by race. This time, instead of pulling the data into Julia with st_matrix(), I copy the entire data set from Stata with jl save. Since I specify no name for the target DataFrame, it defaults to df:
  • jl save
    jl
      dropmissing!(df, :bmi)
      races = ["Other", "White", "Black"]
      f = Figure()  # make new, empty figure
      Axis(f[1, 1], yticks=((1:3)/20, races), xlabel="BMI")  # add empty plot region
      for (i,r) in enumerate(races)
        density!(df.bmi[df.race .== r], offset=i/20, color=(:slategray, .4), bandwidth=1)
      end
      f |> display
    exit()
dropmissing!(df, :bmi) is similar to drop if bmi==. in Stata. However, it changes the type of df.bmi from Float32? to Float32, meaning that the column no longer accepts missing values. That is apparently necessary to make Makie work. The “!” in “dropmissing!” has no special meaning in the syntax of Julia: it is just a character in the function’s name. By convention, functions that modify their arguments have names ending in “!”.
The for...enumerate() loop iterates over the racial groups. r takes race names as string values while i is a counter starting from 1. In the Stata data set, the race column is coded with 1, 2, and 3. Those codes are displayed through a value label as “White”, “Black”, and “Other.” Since jl save stores race as a CategoricalVector, the Julia column’s contents are outwardly strings, not numbers. df.bmi[df.race .== r] uses Boolean indexing to extract the elements of the bmi column for which race takes the string value in r.
Each iteration plots a density in slate gray with 40% opacity. The spacing offset of $1/20$ is interpreted in the vertical axis’s units, BMI density, and was chosen by trial and error to cause overlapping. See Figure 2.
Figure 2: Overlapping density plots made with Makie in Julia
4 Tips on using the jl command
4.1 The JuliaMono font
Julia output routinely contains certain Unicode characters, such as vertical and diagonal ellipses, that Stata fonts do not cover. (However, this is not necessarily the case for Stata packages that call Julia back ends.) The JuliaMono font is a free, monospaced font that includes those characters, and looks reasonably good in Stata’s results window. The font can be installed from github.com/cormullion/juliamono/releases. Type help fonts in Stata to learn how to change the Results window font.
4.2 Multithreading
Julia includes several facilities for multitasking. One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other. reghdfejl’s vce(bs) option uses that feature. A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space. That is the sort @tturbo exploits. The maximum number of threads Julia allows is set when it launches, and it can default to 1. A good guess for the optimum is the number of “performance” or “big” cores on the CPU. You can determine the current limit with the Julia command Threads.nthreads().
How best to set Julia’s thread limit depends on the operating system. In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS. In Linux, you can add an entry such as export JULIA_NUM_THREADS = 6 to $\sim$/.bashrc” or the equivalent. In macOS, environment variables do not affect graphical applications such as Stata when launched from the desktop. One can instead set the limit in each Stata session through the threads() option of the jl start command. Example: jl start, threads(8). This must be the first jl command issued since starting Stata. It must also precede invocation of Julia-dependent programs such as reghdfejl. [^11] In macOS, one can in fact assign JULIA_NUM_THREADS in $\sim$/.zshenv”. This will take effect when one launches Stata from the terminal. For example open -a StataMP will launch Stata/MP.
4.3 Limitations on accessing Stata objects
The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details. However initiated, the job of copying falls to the C++ plugin. To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface. [^12] See stata.com/plugins. The SPI imposes certain constraints. For one, it makes public no functions that can create Stata variables or matrices. As a result, you can edit existing Stata variables and matrices from Julia, but not create new ones. The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment. For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy from a Julia matrix.
A subtler limitation pertains to local macros. Explaining it requires a deeper dive. The “jl.ado” file defines the jl command. In order to call the plugin “jl.plugin,” it first declares the plugin’s existence with this line:
  • program _julia, plugin using(jl.plugin)
You are not supposed to call the plugin directly. But you could. For example:
  • plugin call _julia, eval "1+1"
The designated entry point to the plugin, the C++ function stata_call(), would then receive and process the two arguments after the comma, which tell it to have Julia evaluate “1+1”. Since calls to the plugin are in general arcane and sometimes require preparatory steps, the jl program serves as an essential and more user-friendly wrapper. That extra layer complicates access to local macros. If Julia, having been invoked by jl and the plugin, uses the SPI function SF_macro_save() to write a Stata local named foo, then foo will appear in the macro name space of the program that called the plugin, which is jl. It will not appear in the context a step up, which could be your code. As a workaround, every time the Julia function st_local() is called, it adds the name of the macro it is writing to another local, __jllocals. Before exiting, jl copies every macro in that list to its caller’s context using the undocumented Stata command c_local.
However, the same complication prevents one from reading Stata locals from Julia. The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros. In partial compensation, single-line jl: commands can quote locals, because Stata processes the lines before passing them to jl. For example:
  • local varname x1
    jl: sum(df.`varname')
This trick will not work in an interactive Julia session.
4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
The specification of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double. By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types. [^13] Because the ranges of Stata’s byte, int, and long types are narrowed in the way described in the next paragraph, when transferring data from Julia to Stata, the corresponding Julia types are mapped to larger types in Stata. For example, columns of type Int8, which have an allowed range of $[-128,+127]$, are copied to Stata’s two-byte int type, since Stata bytes only hold $[-128,100]$. So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame. The jl commands can do nothing about the first conversion. But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision. This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive all the data, rather than a separate block for each of variously typed columns.
Separately, Stata and Julia represent missing values in different ways. Stata reserves certain high values of each data type for 27 flavors of missing: . and .a, …, .z. That is why the maximum value of a Stata byte is 100, not than 127. Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values. It also allows observations of string variables to be missing. The julia package translates missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices). For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. The julia plugin contains optimized code that runs when such clauses are not specified.
Data transfers are thus fastest when they map double precision to double precision and include no missing values or sample restrictions. Consider this example, which creates a 10 million $\times$ 10 set of double-precision numbers and copies them to a Julia DataFrame named demo:
  • set obs 100000000
    drawnorm double x1-x10
    jl PutVarsToDF x1-x10, nomissing doubleonly dest(demo)
On my computer, the last line takes 1.3 seconds, for a transfer rate of 6 gigabytes/second.
4.5 Simulated read-eval-print loop
The Julia language system has many components—for interpreting and compiling code, managing memory, and so on. The core parts are written in C++ and distributed as compiled libraries. Other components are themselves written in Julia, including the front end, which users often think of as Julia per se. The front end is usually called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat. But the REPL provides other conveniences: typing ”?” starts help mode; typing ”]” instead switches to the Julia package manager; some use is made of color; arrow keys allow one to explore one’s command history.
While there may be a way to run the full Julia REPL inside of Stata, the jl command does not do that. It gins up a limited REPL. That is because The plugin directly accesses lower-level components of Julia, through the C-language entry points that loosely comprise Julia’s interface for “embedding” in C. A workhorse for jl is the Julia system’s jl_eval_string() function. It accepts a C string holding a Julia expression, evaluates it, and returns the result in a C structure that represents a generic Julia object. All Julia lines submitted to jl pass through this function. As its name suggests, the job of the function is to evaluate strings, not provide a user experience.
Especially when in interactive mode, jl takes steps to create an ersatz Julia REPL. As described in the first example in section 2, it checks whether lines are syntactically complete and prompts the user to continue typing if not. That allows one to enter a multiline code block without triggering a syntax error when it is still incomplete. jl also captures printed output (from stdout) in order to display not only the return value of a command but anything printed while running it. jl translates lines beginning with “?” into help requests and suppresses output from lines ending with “;”.
Another REPL feature that jl simulates matters for code such as this:
  • jl
      S = 0
      for i in 1:10
        S += i
      end
      print(S)
    exit()
This snippet prints the sum of the first 10 whole numbers—at least when run in the Julia REPL. Actually, for reasons having nothing to do with Stata, the snippet will crash if run non-interactively, as part of a stored program. Under Julia’s “soft scoping” rule, since S is not declared or initialized within the for loop before the first attempt to increment it, the attempt will trigger an undefined variable error. Because that rule is counterintuitive, Julia’s REPL engineers an exception for interactive use. [^14] See the documentation at docs.julialang.org/en/v1/manual/variables-and-scoping/#local-scope. jl: imitates the exception by calling the Julia softscope() function on the user’s input after parsing it. softscope() is itself part of the code base of the Julia REPL.
Still, jl’s REPL simulation is incomplete. Typing ? and ] alone have no effect. Plots do not automatically appear, which is why the examples in section 3.3 explicitly call display().
And the processing required for the simulation, along with the special handling of Stata locals, adds time to the execution of each Julia line from Stata. The extra 0.02 seconds or so is not noticeable in an interactive session. But it can add up in a program that calls jl: many times. The _jl: variant of the prefix command bypasses all of this processing.
5 The julia package
5.1 Architecture
The julia package consists of three intertwined pieces, all of which have been mentioned. The core is a shared library written in C++ according to the specification for Stata plugins. It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending commands to Julia and returning the results.
The file “jl.ado” serves as a Stata front end for the plugin, providing a more intuitive interface and performing tasks better done in the ado language than in C++. At the suggestion of Jeffrey Pitblado at StataCorp, this program is called jl rather than julia because Stata might introduce an official julia command someday. When jl is first called in a Stata session, it briefly runs the Julia REPL through a shell command in order to query the location on disk of Julia’s shared libraries. It passes the location to the plugin, which loads the libraries. This lets the plugin access Julia.
jl also calls Stata’s findfile to track down the third key piece of the package. “stataplugininterface.jl” defines a module of Stata interface functions in Julia, such as st_local() and st_data(). These work by calling entry points in the Stata plugin interface, which are exported to Julia by the plugin. (The Julia language includes a way to call C functions.) At start-up, jl therefore stores the location of “jl.plugin” in the stataplugininterface module.
5.2 Stata commands in the julia package
As we have seen, jl: and the faster _jl: prefix commands run lines of Julia code:
  • jl: juliaexpr _jl: juliaexpr
where juliaexpr is an expression to be evaluated in Julia.
jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to set the upper limit on the number of threads:
  • jl start [, threads(# | auto)]
Specifying threads(auto) instructs Julia to use its own heuristic to pick the number of threads. Executing jl start is optional, for jl will initialize itself as needed.
Several subcommands help manage Julia packages. Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments. Julia also makes it easy to change the current environment. This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration. The package management subcommands are:
  • jl GetEnv
    jl SetEnv [name]
    jl AddPkg name, [minver(X.Y.Z)]
GetEnv displays the directory of the current package environment and returns it in r() macros. On my computer, which is running Julia 1.10, the command displays:
  • . jl GetEnv
[…]
jl SetEnv moves to a chosen, named environment, which is merely a subdirectory of the default environment directory:
  • . jl SetEnv MyEnv
[…]
If the environment does not exist, it is created and populated with the DataFrames.jl and CategoricalArrays.jl packages. One can return to the default environment with SetEnv . or just SetEnv. [^15] In distributed computing, Julia worker processes do not automatically inherit the master’s package environment. One can pass on the master’s environment through command line switches when launching the workers, such as in addprocs(4, exeflags="--project=\$(Base.active_project())").
Last in the package management family is the AddPkg option, which installs a package in the current environment or, if minver() is specified, updates it to the latest version if the current version falls below the minimum. Versions X.Y.Z are in the three-part semantic versioning format.
The remaining jl subcommands copy data between Stata and Julia. The first six listed below are meant more as programmer’s commands while the last two are higher-level. jl GetVarsFromDF, for example, requires that the Stata data set already be at least as tall as the DataFrame from which it will receive variables. In contrast, jl use works like Stata’s use command: it replaces the current data set, and will fail unless the replace option is included, or the current data set has not changed since it was last saved.
  • jl PutVarsToDF [varlist] [if] [in], [destination(string) cols(string) nolabel nomissing doubleonly] jl GetVarsFromDF varlist [if] [in], [cols(string) source(string) replace nomissing] jl PutVarsToMat [varlist] [if] [in], destination(string) jl GetVarsFromMat varlist [if] [in], source(string) jl PutMatToMat matname, [destination(string)] jl GetMatFromMat matname, [source(string)] jl save [dataframename], [nolabel nomissing doubleonly] jl use dataframename, [clear] jl use varlist using dataframename, [clear]
The varlists and matnames before the commas refer to Stata variables or matrices. If an optional varlist is omitted, it defaults to *, meaning all variables in the current data set in their current order. A destination() or source() option after the comma refers to a Julia matrix or DataFrame. When an optional DataFrame name is not provided, it defaults to df. The cols() option specifies the DataFrame columns to be copied to or from. It defaults to the Stata varlist before the comma. Destination Stata matrices and Julia matrices and DataFrames are entirely replaced. The low-level commands will create the destination Stata variables or, if replace is specified, overwrite them subject to any [if] or [in] restriction.
5.3 Julia functions in the julia package
Functions in the stataplugininterface module can also be characterized as low- or high-level. The low-level ones wrap and mimic functions in the Stata plugin interface. For example, SF_vdata() extracts a single observation on a single variable. The higher-level ones build on the low-level ones in order to imitate some of Mata’s Stata interface functions. st_view(), for instance, returns a Julia view onto one or more numeric Stata variables, optionally restricting to a subsample. Through the view, Julia code can read and write Stata variables while treating them as part of a matrix.
The full list of functions:
SF_nobs() Number of observations in Stata data set
SF_nvar() Number of variables in Stata data set
SF_var_is_string(i) Whether variable i is string
SF_var_is_strl(i) Whether variable i is a strL
SF_var_is_binary(i, j) Whether observation i of variable j is a binary strL
SF_sdatalen(j, i) String length of variable i, observation j
SF_is_missing() Whether a Float64 value is Stata missing
SV_missval() Stata floating-point value for missing
SF_vstore(j, i, val) Set observation j of variable i to val (numeric)
SF_sstore(j, i, s) Set observation j of variable i to s (string)
SF_vdata(j, i) Get observation j of variable i (numeric)
SF_sdata(j, i) Get observation j of variable i (string)
SF_macro_save(mac, tosave) Set macro mac
SF_macro_use(mac, maxlen) First maxlen characters of macro mac
SF_scal_save(scal, val) Set value of scalar scal
SF_scal_use(scal) Get scalar scal
SF_row(mat) Number of rows of matrix mat
SF_col(mat) Number of columns of matrix mat
SF_mat_store(mat, i, j, val) mat[i,j] = val
SF_mat_el(mat, i, j) Get mat[i,j]
SF_display(s) Print to Stata results window
SF_error(s) Print error to Stata results window
st_nobs() Same as SF_nobs()
st_nvar() Same as SF_nvar()
st_varindex(s) Index of variable named s in data set
st_global(mac) Get global macro mac
st_global(mac, tosave) Set global macro mac
st_local(mac, tosave) Set local macro mac
st_numscalar(scal) Get scalar scal; same as SF_scal_use()
st_numscalar(scal, val) Set scalar scal; same as SF_scal_save()
st_matrix(matname) Get numeric Stata matrix
st_matrix(matname, jlmat) Put Julia matrix in pre-existing Stata matrix
st_data(varnames) Get Stata variables, as matrix
st_data(varnames, sample) Get Stata variables, with sample restriction, as matrix
st_view(varnames) Get Stata variables, as view
st_view(varnames, sample) Get Stata variables, with sample restriction, as view
The limitations noted in section 4.3 can be seen in the list. While st_global() can read and write globals, st_local() can only write locals. st_matrix() can only write to Stata matrices that already exist.
An interactive session demonstrates the use of some of the functions:
  • sysuse auto
    jl
      sample = rand(Bool, st_nobs())    # random Boolean vector defining subsample
      v = st_view("price mpg", sample)  # view onto subsample of two vars
      v ./= 2                           # halve these Stata data points
      st_numscalar("s", sum(v))         # sum the halved data into Stata scalar s
      st_local("m", string(st_numscalar("s"))) # put sum as string in Stata local m
    exit()
One can obtain somewhat more information about each function through the help facility, for example with “jl:?st_view”.
6 Conclusions
Julia is a strong option for universal, back-end development of statistical applications. Working in Julia comes with some drawbacks, including the challenge of learning a complex new language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation. But Julia and nearly all Julia packages are free and open and run on all popular operating systems. The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain. The package ecosystem covers domains essential to efficient statistical work, such as multiprocessing, linear algebra, and numerical optimization.
Instead of implementing new methods separately in Stata or R or Python, or in a low-level language such as C, developers can write the back ends once, in Julia. Here, WildBootTests.jl shows the way: it is an optional back end for boottest in Stata and a required back end for instrumental variables inference with fwildclusterboot in R.
For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia. Through a plugin, data can be piped at high speed. Most features needed to make Stata objects accessible in Julia can and have been implemented. The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed, new features, or more flexible syntax (as for plotting). reghdfejl is a perhaps the “killer app” for the Julia-Stata link. It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for reghdfe. And reghdfejl preliminarily handles nonlinear HDFE models, which brings new functionality to Stata users.
StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java. Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables. They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols. They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it first. They could make Mata interoperate with Julia.
The goal for the julia package was to prove a concept. In that I think it has succeeded.

### v3

@@
 Abstract
-Like Python and Java, which are integrated into Stata, Julia is a free programming language that runs on all major operating systems.
-The julia package links Stata to Julia as well.
-Users can transfer data between Stata and Julia at high speed, issue Julia commands from Stata to analyze and plot, and pass results back to Stata.
-Julia’s econometric software ecosystem is not as mature as Stata’s or R’s, or even Python’s.
+The julia package integrates the Julia programming language into Stata.
+Users can transfer data between Stata and Julia, issue Julia commands to analyze and plot, and pass results back to Stata.
+Julia’s econometric ecosystem is not as mature as Stata’s or R’s or Python’s.
 But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms.
-The boottest program for wild bootstrap–based inference (Roodman et al., 2019) can call a Julia back end for a 33–50% speed-up, even as the R package fwildclusterboot (Fischer and Roodman, 2021) uses the same back end for inference after instrumental variables estimation.
-reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls an independently developed Julia package for tenfold acceleration on hard problems.
-reghdfejl also supports nonlinear models—preliminarily, as the Julia package for that purpose matures.
-NOTE: This article is formatted for submission to the Stata Journal. It has not been accepted.
+For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can both call the same Julia back end.
+And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls a Julia package for tenfold acceleration on hard problems.
+reghdfejl also supports nonlinear fixed-effect models that cannot otherwise be fit in Stata—though preliminarily, as the Julia package for that purpose is immature.
 keywords:
@@
 In Stata, for example, a call to regress or gsem starts out in the ado programming language.
-The ado code translates your command line into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language.
+The ado code translates your command into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language.
 In modern Stata, an ado program can even call Python or Java.
@@
 Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution.
-(Python and C++ is a common combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that code for reuse.
-Like LISP, a language almost as old as Fortran, Julia can treat code as data.
-This feature opens the door to meta-programs that can, for example, automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
+(Python and C++ is a common two-language combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse.
+A final strength is that, like LISP, Julia can treat code as data.
+This feature opens the door to meta-programs that can automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
 Julia has disadvantages.
 While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented.
-Standards for passing results are less well developed than in Stata.
-Much less statistical functionality is available in Julia than in Stata or R, or even Python.
+Standards for storing and returning results are less well developed than in Stata.
+Much less statistical functionality is available than in Stata or R, or even Python.
 Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development.
 Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms.
-The core work can implemented, optimized, and augmented in one place, rather than separately on each platform.
+The core work can augmented and optimized in one place, rather than separately on each platform.
 The julia package links Stata to Julia.
-It follows in the footsteps of many projects to link statistical software to other languages.
+It follows in the footsteps of projects to link statistical software to other languages.
 StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003.
-It added the built-in Mata language in Stata 9.
+It added Mata in Stata 9.
 Fiedler (2012) envisioned running Python inside of Stata.
-Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a plugin for Stata.
+Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a C++ plugin for Stata.
 Perhaps inspired by Fiedler, Stata 16 fully integrated Python.
@@
 Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk.
-Outside of Stata, the JuliaCall package (Li, 2019) lets R users call Julia routines.
-An identically named package does the same for Python, and is paired with PyCall for Julia.
+Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (JuliaConnectoR) packages link R to Julia.
+Another package named JuliaCall does the same for Python.
 [^1] See juliapy.github.io/PythonCall.jl/stable.
 The julia package was most inspired by Fiedler (2013).
-Its core too is a C++ plugin, which comes precompiled for Windows, Linux, macOS/Intel, and macOS/ARM.
+Its core too is a C++ plugin, which comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon.
 From the user’s point of view, the package has three components:
@@
   • Julia functions to read and write Stata macros, scalars, matrices, and variables.
-The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package.
+The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package for Stata.
 Two user-written Stata packages call Julia via julia.
-The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata back end.
-And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016), while calling FixedEffectModels.jl, whose lead developer is Matthieu Gomez.
+The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata implementation I originally wrote for boottest.
+And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016).
+reghdfejl calls the Julia package FixedEffectModels.jl, whose lead developer is Matthieu Gomez.
 reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl.
@@
 Still, this paper explains the new julia package.
-So it starts with a hands-on demonstration of running Julia from Stata, in section 2.
-Then section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia.
+It starts by demonstrating how to run Julia from Stata, in section 2.
+Section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia.
 Section 4 gives tips and warns of some limitations.
@@
 The preferred way to get Julia is to download the juliaup installation manager.
-You do that in Windows by installing “Julia” (not “juliaup”) from the Windows Store.
+You do that in Windows by installing Julia (not juliaup) from the Windows Store.
 [^3] The instructions might change.
@@
     "Hello world!"
-The output may take several minutes to appear.
-On first use, jl installs and precompiles two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on.
-DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets—i.e., rectangular grids whose columns have names and potentially different data types.
+This operation may take several minutes.
+In Windows, many mysterious console windows may appear and disappear as required libraries are downloaded and pre-compiled.
+In macOS and Linux, output from the installation process may appear in terminal windows.
+For on first use, jl installs two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on.
+DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, i.e., rectangular grids whose columns have names and potentially different data types.
 CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
 In general when using Julia, you will experience delays at four points: when packages are installed; when they are first used; when they are first used in a Julia session; and when particular features within them are first used, such as clustering standard errors.
-A thrust of recent development in the Julia langauge has been to move the behind-the-scenes processing from later to earlier stages in order to shorten the lags during regular use.
-In the session shown below, I load Stata’s venerable “auto” data set, extract five variables, run a regression on them, and then copy the data to a Julia DataFrame named auto:
+A thrust of recent development in the Julia language has been to move the behind-the-scenes processing to earlier stages in order to shorten the lags during regular use.
+In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them, then copy the data to a Julia DataFrame named auto:
   • . sysuse auto
@@
     Data saved to DataFrame auto in Julia
-jl save is not a Julia command. It is a subcommand of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten.
-Next I type jl by itself, which changes the prompt in much the same way that typing mata or python or java does.
-Have started an interactive Julia session, I display the new DataFrame by typing its name:
+jl save is not a Julia command.
+It is a subcommand of the Stata command jl, which is part of the julia package.
+If an object named auto already exists in the Julia environment, it is overwritten by the incoming copy of the Stata variables.
+Next I type jl by itself, which changes the prompt the same way that typing mata or python or java does.
+It starts an interactive Julia session, in which I can type multiple Julia commands.
+I start with displaying the new Julia DataFrame by typing its name:
   • . jl
@@
 jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector.
-While the Julia version of the variable also has an internal, integer representation, the vector presents as a text variable.
+While this Julia version of the variable is also stored as a column of integers—1’s and 2’s—it presents as a text variable.
 [^4] Julia categorical variables are not quite comparable to Stata variables with data labels.
-Code involving the Julia variable refers to display values, such as “Domestic,” not to numerical codes.
-The type label “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
-Next I run the same regression in Julia, with the reg() function from the Julia package FixedEffectModels.jl:
-  • jl> . using FixedEffectModels
+Commands involving the Julia variable refer to display values such as “Domestic,” not to numeric codes.
+The type name “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
+Next I run the same regression as before, but now in Julia, with the reg() function of the Julia package FixedEffectModels.jl:
+  • jl> . jl AddPkg FixedEffectModels  # assure FixedffectModels installed
+    jl> . using FixedEffectModels      # load it for use
     jl> . reg(auto, term(:price) ~ term(:headroom) + term(:foreign)&term(:mpg),
-      ...     cov.cluster(:turn))
+      ...     Vcov.cluster(:turn))
     FixedEffectModel
@@
 After I hit the enter key at the end of the first line, jl detected that the line was not grammatical by itself and printed “...” to prompt me to continue.
-  • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears here.
-  • All variable names in the command line are prefixed with “:”. That indicates that the names are symbols, not stand-alone Julia objects like the DataFrame auto.
-  • Because foreign is a CategoricalVector, reg() automatically treats it as a factor variable, generating a pair of dummies to be interacted with mpg.
-There is no need for an “i.” prefix or the equivalent as in Stata.
-However, treating non-CategoricalVectors as factor variables requires a more awkward syntax.
+  • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears above.
+  • All variable names in the command line are prefixed with “:”.
+That indicates that the names are symbols, not stand-alone Julia objects like the DataFrame auto.
+The symbols are names of columns in auto.
+  • Because foreign is stored as a CategoricalVector, reg() automatically treats it as a factor variable, generating a pair of dummies to be interacted with mpg.
+There is no need for a Julia equivalent of Stata’s “i.” prefix.
+However, in Julia, treating non-CategoricalVectors as factor variables requires a more awkward syntax.
 One must pass reg() an argument with a dictionary mapping the variables to coding types, such as with contrasts=Dict(:headroom=>DummyCoding()).
@@
 In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model.
-The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and surfaces for the user.
+The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and makes available to the user.
 The content of each Term is just a symbol, such as :price.
 What give Term’s life are the operations that StatsModels.jl defines on them.
-The + operator is repurposed (“overloaded”) to form tuples of Term’s—pairs, triplets, and so on.
-The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides.
-The entire FormulaTerm is passed as the second argument to reg().
+The + operator is repurposed (“overloaded”) to form lists (tuples) of Term’s.
+The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides, just as in a linear regression model.
+In the example, an entire FormulaTerm is passed as the second argument to reg().
 In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
@@
 The StatsModels.jl formula language is also extensible [^5] See juliastats.org/StatsModels.jl/stable/internals/#extending..
-As we will see, FixedEffectModels.jl defines a way to mark variables whose fixed effects are to be absorbed.
+As we will see, FixedEffectModels.jl extends the formula language in order to mark variables whose fixed effects are to be absorbed.
 Yet a powerful feature of Julia makes it practical to ignore most of this complexity.
 The regression command above can be simplified with a macro.
-Where macros in Stata are strings, in Julia they are code for manipulating code—meta-programs.
-StatsModels.jl defines @formula, which lets one write models in a more R-like way.
+Where macros in Stata are strings, in Julia they are code for manipulating code: meta-programs.
+StatsModels.jl defines a macro called @formula, which lets one write models more simply.
 Thus:
@@
     [identical output omitted]
-Before this new command is compiled and executed, @formula turns it into something closer to the previous one.
+Before this new command is compiled and executed, @formula turns it into something closer to the earlier regression command.
 [^6] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones.
 An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
-Notice one other change to the command line: it now starts with m= in order to store the fitted model in the Julia variable m. This lets me extract return values and pass them back to Stata:
+Notice one other change to the regression command: it now starts with m = in order to store the fitted model in the Julia variable m. This lets me extract return values and pass them back to Stata:
   • jl> . st_numscalar("adjR2", adjr2(m))
@@
 2.2 Built for speed
-To partially convey why Julia is a good environment for developing numerical software, I present another example.
+To partially convey why Julia is a good environment for scientific computing, I present another example.
 Suppose we have a data set with 10 million rows and 10 columns named x1-x10.
@@
           N, M = size(X);                              ///
-          retval = zeros(N,1);                         ///
+          retval = zeros(N);                           ///
           @tturbo for i in 1:N                         ///
@@
 To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line.
-It is typographically split across multiple lines with Stata’s continuation token, ///.
-When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon.
+It is typographically, not syntactically, split across multiple lines with Stata’s continuation token, ///.
+When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
 The code for this new Julia function, XQX(), is mostly straightforward.
@@
 The macro may even reorder the nesting of the loops with an eye toward the fact that in Julia, matrices are stored column by column, and on-chip memory caching makes it is faster to access adjacent memory locations in sequence.
-The LoopVectorization package is called that because it also tailors code so that the resulting machine language exploits the ability of modern chips to vectorize operations within each core, in 256- or 512-bit registers—what is called single instruction, multiple data (SIMD) execution.
+It may tailor the code so that the resulting machine language exploits the ability of modern chips to vectorize operations within each CPU core, in 256- or 512-bit registers—what is called single instruction, multiple data (SIMD) execution.
 And @tturbo will exploit multithreading, the subdivision of work across cores; that is what the first “t” in tturbo stands for.
 Triply nested implementations in Mata and Python are much slower, I assume because the calculations in the innermost loop have to be interpreted a billion times.
-The best alternatives I have found in the two languages are essentially the same:
+In constrast, Julia translates all the code down to machine language just once.
+The best alternatives I have found in the other two languages are essentially the same:
   • mata
@@
 However, both contain a subtle inefficiency.
-They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$.
-Since in our example ${\bf X}$ is $10^{6}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each.
+They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$, before summing their rows and dispensing with them.
+Since in our example ${\bf X}$ is $10^{7}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each.
 Allocating and deallocating memory itself takes time.
 Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory.
-The triply-looped Julia code demands essentially no temporary storage.
+In contrast, the triply-looped Julia code demands essentially no temporary storage.
+This inefficiency of Mata and Python can be seen as an artifact of the two-language problem Julia was designed to solve.
+The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are nearly impossible to revise, yet not optimized for the purpose at hand.
 To create a test bed for these functions, I construct ${\bf X}$ as a Stata data set and ${\bf Q}$ as a Stata matrix, then copy them into Mata, Python, and Julia.
-The set rmsg on command generates timings:
+The set rmsg on command generates timing reports:
   • . python: import numpy as np
@@
     r; t=0.15
-Under a Julia convention, the semicolons at the end of the jl lines suppress the printing of the results of the assignments.
-The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because of just-in-time compilation.
-On a Windows laptop with an Intel i9-13900H, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python.
+Under a Julia convention, the semicolons at the end of the jl lines suppress printing of the results of the assignments.
+The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because that is when it is compiled down to machine code.
+On a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python.
 In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
 To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above.
-In order to trigger immediate compilation, the Julia package manager would run the script after installing the package.
+The Julia package manager would run the script after installing the package in order to trigger immediate compilation.
 Julia would save the generated machine code for future use.
@@
 Julia packages for HDFE modeling are also available: Matthieu Gomez’s FixedEffectModels.jl and Johannes Boehm’s GLFixedEffectModels.jl.
-With regard to linear models, according to benchmarking results on the GitHub pages of fixest and FixedEffectModels.jl, those two packages are 1–2 orders of magnitude faster than regdhfe.
+According to benchmarking results on the GitHub pages of fixest and FixedEffectModels.jl, those two packages are 1–2 orders of magnitude faster than regdhfe.
 [^7] See github.com/lrberge/fixest#benchmarking and github.com/FixedEffects/FixedEffectModels.jl#benchmarks.
-These developments have given R and Julia users faster and more general implementations of HDFE modeling than were available in Stata.
+These developments have given R and Julia users faster and more general implementations of HDFE modeling than have been available in Stata.
 That is what motivated the development of the julia and reghdfejl packages.
-The idea was to not reinvent wheels, but to fashion a familiar-looking chassis onto existing wheels.
+The idea was to not reinvent wheels, but to fashion a familiar-looking Stata chassis onto existing Julia wheels.
 The extended example in section 2.1 followed a sequence that is the essence of reghdfejl: copy data from Stata to Julia, estimate in Julia, return results to Stata.
@@
 In practice, the two differ outwardly as well as inwardly.
-As reghdfjle depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience.
+As reghdfjl depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience.
 reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering.
-So it often reports slightly larger standard errors.
+So it sometimes reports slightly larger standard errors.
 While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007).
 As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic.
-reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and pass the results to ivreg2.
+reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and then transfer control to ivreg2.
 But in this mode, reghdfejl is no faster than ivreghdfe.
-For on hard problems, ivreg2 dominates the runtime.
+For on hard problems, ivreg2 accounts for most of the run time.
 reghdfe and reghdfejl share the following options: absorb(), vce(), residuals(), tolerance(), iterate(#), nosample, keepsingletons, compact, level(#), and standard Stata regression display options.
@@
 For instance, a(firm year_fe=year) will absorb firm and year fixed effects, and store estimates of the latter in the variable year_fe.
-For both reghdfe and reghdfejl, the tolerance() option sets a precision threshold for the fitting algorithms to declare convergence; however, the semantics are not necessarily the same, so that, say, tol(1e-6) could elicit slightly different results from the two.
-reghdfejl accepts three distinctive (documented) options:
+For both reghdfe and reghdfejl, the tolerance() option sets a precision threshold for the fitting algorithms to declare convergence; however, the semantics are not necessarily the same, so that, say, tol(1e-6) could elicit slightly different results from the two programs.
+reghdfejl accepts three distinctive options:
 gpu requests the use of a GPU. Currently the option saves more time on NVIDIA than Apple GPUs.
-threads(#) lowers the limit on the number of threads that can be used. It cannot increase the thread count above the limit set when Julia is started (see section 4.2). This option is rarely used.
-verbose directs reghdfejl to show more of its work—to display the Julia copy of the data set, the formula for the regression model, and the regression command.
-The data set and formula are left behind for the user to work with through jl, not erased as they otherwise would be.
+threads(#) caps the number of parallel threads that can be used. It cannot increase the thread count above the limit set when Julia is started (see section 4.2). This option is rarely used.
+verbose directs reghdfejl to show more of its work—to display the Julia copy of the data set, the formula for the regression model, and the Julia regression command.
+The data set and formula are left behind for the user to work with  not erased as they otherwise would be.
 In addition, reghdfejl offers a bootstrap variance-covariance estimator. vce(bootstrap) (or vce(bs)) is standard within Stata; regress accepts it. As an example, for reghdfejl, the clause
   • vce(bs, reps(1000) seed(42) procs(4) cluster(year))
-specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from clusters defined by year.
+specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from groups defined by year.
 One could attain the same results with bootstrap:reghdfe or bs:reghdfe.
 But reghdfejl implements vce(bs) in a radically faster way.
-After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times, etc.).
-To compare reghdfe and reghdfejl on speed, I create a data set with 10 million rows and two sets of fixed effects.
-I run the same regression with both commands.
-I explore the effect of switching from one to multiple processors in Stata/MP, and of adding reghdfejl’s gpu option:
+After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times in a bootstrap sample, and so on.).
+To demonstrate reghdfejl’s speed, I create a data set with 10 million rows and two sets of fixed effects.
+I run the same regression with reghdfe and reghdfejl, as well as with the built-in areg command, which in StataNow 18.5 gained the ability to absorb multiple fixed effects.
+While reghdfe and reghdfejl can cluster standard errors on multiple variables (CGM_2011), areg cannot, so the example is restricted to one-way clustering:
   • set obs 10000000
@@
     set processors 1
-    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
-    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
-    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
+    areg      y x1 x2, a(id1 id2) cluster(id1)
+    reghdfe   y x1 x2, a(id1 id2) cluster(id1) dof(none)
+    reghdfejl y x1 x2, a(id1 id2) cluster(id1)
+    reghdfejl y x1 x2, a(id1 id2) cluster(id1) gpu
     set processors 6
-    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
-    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
-    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
+    areg      y x1 x2, a(id1 id2) cluster(id1)
+    reghdfe   y x1 x2, a(id1 id2) cluster(id1) dof(none)
+    reghdfejl y x1 x2, a(id1 id2) cluster(id1)
+    reghdfejl y x1 x2, a(id1 id2) cluster(id1) gpu
 For comparability, the reghdfe lines include dof(none) to disable the degrees-of-freedom correction, since reghdfejl lacks that feature.
-The first triplet runs on one CPU core, as under a non-MP flavor of Stata.
-But that setting on restricts the ado and Mata code: the Julia code still runs on multiple CPU threads.
-On a laptop with an Intel i9-13900H CPU and an NVIDIA RTX 4070 GPU, reghdfe takes 32 seconds.
-reghdfejl takes 3.8 seconds without the gpu option and 2.8 with.
-Moving to 6 processors in Stata and Mata—since the Intel chip has 6 “performance” cores—lowers the times to 24, 3.3, and 2.7 seconds.
+The first quadruplet runs on one CPU core, as under a non-MP flavor of Stata.
+But that setting only affects the execution of reghdfe’s ado and Mata code: the Julia program still runs on multiple CPU threads.
+On a laptop with an Intel i9-13900H CPU and an NVIDIA RTX 4070 GPU, areg takes 8.1 seconds and reghdfe 17 seconds.
+reghdfejl takes 2.6 seconds without the gpu option and 1.8 with.
+Moving to 6 processors in Stata and Mata—since the Intel chip has 6 “performance” cores—lowers the times to 5.0, 10, 2.2, and 1.8 seconds.
 Another script benchmarks reghdfejl’s vce(bs) option:
@@
       reghdfe ln_wage age ttl_exp tenure not_smsa south, absorb(year occ_code) dof(none)
-    parallel initialize 6
+    parallel initialize 12
     parallel bs, cluster(occ_code) reps(1000): ///
@@
       vce(bs, cluster(occ_code) reps(1000) procs(12))
-Using Stata’s bs prefix command with reghdfe takes 64 seconds.
-Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 6 copies of Stata [^8] See github.com/gvegayon/parallel.
-cuts the time to 23 seconds.
-reghdfejl takes just 2.5 seconds.
-But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 6 worker copies of Julia and triggers some just-in-time compilation in each.
-The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers is kept at 6.
-The reghdfejl package includes a partialhdfejl command for absorbing fixed effects out of other variables.
+Using Stata’s bs prefix command with reghdfe takes 69 seconds.
+Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 12 copies of Stata [^8] See github.com/gvegayon/parallel.
+cuts the time to 19 seconds.
+reghdfejl takes just 1.9 seconds.
+But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 12 worker copies of Julia and triggers some just-in-time compilation in each.
+The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers reqested by the procs() suboption of the vce() option does not rise further.
+The reghdfejl package includes a partialhdfejl command for absorbing (partialling) fixed effects out of other variables.
 partialhdfejl command lines must contain either a generate() option to name the variables that will hold the results, or a prefix() option for making the new variable names out of the old.
@@
 (2019) introduces the Stata package boottest for inference via the wild bootstrap.
-Among the methods boottest implements is the “wild restricted efficient” bootstrap (WRE), which is the Davidson and MacKinnon (2010) adaptation for instrumental variables (IV) estimation.
+Among the methods boottest implements is the “wild restricted efficient” bootstrap (WRE), which is the Davidson and MacKinnon (2010) adaptation of the wild bootstrap for instrumental variables (IV) estimation.
 Since the publication of Roodman et al.
@@
 The partialling-out accelerates the application of the user’s IV estimator despite its irreducible nonlinearity (Roodman et al., 2019, section 6.1).
-Also incorporated (more fully), is MacKinnon (2023)’s point that the data only enter wild-bootstrapped inferences through indexed sets of matrix products such as $\{{\bf X}^{\prime}_{g}{\bf Y}_{g}:g=1,...,G\}$ where ${\bf X}_{g}$ and ${\bf Y}_{g}$ hold the rows of data matrices ${\bf X}$ and ${\bf Y}$ for cluster $g$.
+Also incorporated (more fully), is MacKinnon (2023)’s point that the data only enter wild bootstrap computations through indexed sets of matrix products such as $\{{\bf X}^{\prime}_{g}{\bf Y}_{g}:g=1,...,G\}$ where ${\bf X}_{g}$ and ${\bf Y}_{g}$ hold the rows of data matrices ${\bf X}$ and ${\bf Y}$ for cluster $g$.
 Reducing the data to these products early in the calculation complicates the algebra the code must implement, but speeds execution when clusters are few.
-Finally, the back-end functionality in boottest has been transplanted to the Julia package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end.
+And the functionality in boottest, so revised, has been transplanted to the Julia package WildBootTests.jl.
+Adding the julia option to a boottest command line switches to the new back end.
+[^9] Jan Ditzen suggested this interface for calling the Julia implementation.
 Section 8.4 of Roodman et al.
@@
 [^10] Data and preparation code are available at davidroodman.com/david/Levitt.zip.
-Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for adding the julia option to the boottest command line:
+Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for calling boottest to perform inference about changes in the size of the prisoner population affect crime rates:
   • use Levitt, clear
@@
 (2019), needs 210 seconds to perform the two boottest calls.
-The latest version finishes in 7.4 seconds.
-Adding the julia option to the boottest line brings the time to 4.1 seconds.
-The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE.
+The latest version finishes in 7.4 seconds, even without calling Julia.
+Appending julia to the boottest line above lowers the time to 4.1 seconds.
+(However, the first time boottest is called after installation, there is a significant delay for pre-compilation, just as when jl is first run, as described in section section 2.1.)
+The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE, and optionally does so to perform the wild bootstrap after ordinary least squares.
 This may be the first instance of a Julia statistical package serving as a back end on multiple platforms.
 3.3 Plotting
-Many data visualization packages are available for Julia. Surveying them is beyond the scope of this article. To indicate their potential for Stata users, I generate two graphs with Makie.jl.
-Makie works with several plotting back ends, which render output in formats that are optimized for different applications, such as interactive graphs in web pages and publication-quality image files.
+Many data visualization packages are available for Julia.
+Surveying them is beyond the scope of this article.
+To indicate their potential for Stata users, I generate two graphs with the plotting package Makie.jl.
+Makie works with several plotting back ends, which render output in formats designed for different applications, such as interactive graphs in web pages and publication-quality image files.
 The Cairo back end is best for the latter.
 It is bundled with Makie in CairoMakie.jl:
-  • jl AddPkg CairoMakie  // one-time installation
-    jl: using CairoMakie   // use now
-In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.101010blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command.
+  • jl AddPkg CairoMakie  # one-time installation
+    jl: using CairoMakie  # use now
+In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.111111blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command.
 margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights.
 The probabilities are copied to Python and depicted in a three-dimensional surface plot.
-A Julia adaptation of that example is:
+An adaptation of that example to Julia:
   • webuse nhanes2
@@
 This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines.
-The first Julia line imports the data as a matrix and then converts it to a DataFrame with named columns.
+The first Julia line imports the data as a matrix, converts it to a DataFrame, and names the columns.
 Then the surface plot is built and stored in the Julia variable f.
@@
 The next command displays the figure in a new window, which might pop up behind the Stata window.
-That line takes advantage of Julia’s piping syntax for function calls.
-Last, the image is saved to disk.
+That line takes advantage of Julia’s piping syntax for function calls, using the symbol “|>”.
+(display(f) would be equivalent.) Last, the image is saved to disk.
 It appears in Figure 1.
@@
 in Stata.
-However, it changes the type of df.bmi from Float32?
+However, it also changes the type of df.bmi from Float32?
 to Float32, meaning that the column no longer accepts missing values.
@@
 The “!” in “dropmissing!” has no special meaning in the syntax of Julia: it is just a character in the function’s name.
-By convention, functions that modify their arguments have names ending in “!”.
+But by convention, functions that modify their arguments have names ending in “!”.
 The for...enumerate() loop iterates over the racial groups.
-r takes race names as string values while i is a counter starting from 1.
+r takes race names as strings while i is a counter starting from 1.
 In the Stata data set, the race column is coded with 1, 2, and 3.
@@
 df.bmi[df.race .== r] uses Boolean indexing to extract the elements of the bmi column for which race takes the string value in r.
-Each iteration plots a density in slate gray with 40% opacity.
-The spacing offset of $1/20$ is interpreted in the vertical axis’s units, BMI density, and was chosen by trial and error to cause overlapping.
+Each iteration of the loop plots a density in slate gray with 40% opacity.
+The vertical spacing offset of $1/20$ is interpreted in the vertical axis’s units, BMI density, and was chosen by trial and error to cause overlapping.
 See Figure 2.
@@
 (However, this is not necessarily the case for Stata packages that call Julia back ends.) The JuliaMono font is a free, monospaced font that includes those characters, and looks reasonably good in Stata’s results window.
+Without it, some Julia output will appear slightly garbled.
 The font can be installed from github.com/cormullion/juliamono/releases.
@@
 A good guess for the optimum is the number of “performance” or “big” cores on the CPU.
-You can determine the current limit with the Julia command Threads.nthreads().
+You can determine the current limit with the command jl:Threads.nthreads().
 How best to set Julia’s thread limit depends on the operating system.
@@
 Example: jl start, threads(8).
-This must be the first jl command issued since starting Stata.
+This must be the first jl command issued after starting Stata.
 It must also precede invocation of Julia-dependent programs such as reghdfejl.
@@
 For example open -a StataMP will launch Stata/MP.
-4.3 Limitations on accessing Stata objects
+4.3 Limitations in accessing Stata objects
 The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details.
-However initiated, the job of copying falls to the C++ plugin.
+However initiated, the job of copying falls to the C++ plugin included in the julia package.
 To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface.
@@
 The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment.
-For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy from a Julia matrix.
+For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from a Julia matrix.
 A subtler limitation pertains to local macros.
-Explaining it requires a deeper dive.
-The “jl.ado” file defines the jl command.
-In order to call the plugin “jl.plugin,” it first declares the plugin’s existence with this line:
+Explaining it requires a deep dive.
+The “jl.ado” file, also part of the julia package, defines the jl command that is invoked repeatedly in the above examples.
+In order to call the C++plugin, it first declares the plugin’s existence with this line:
   • program _julia, plugin using(jl.plugin)
-You are not supposed to call the plugin directly. But you could. For example:
-  • plugin call _julia, eval "1+1"
-The designated entry point to the plugin, the C++ function stata_call(), would then receive and process the two arguments after the comma, which tell it to have Julia evaluate “1+1”.
-Since calls to the plugin are in general arcane and sometimes require preparatory steps, the jl program serves as an essential and more user-friendly wrapper.
+You are not supposed to call the plugin directly. But after typing that line, you could. For example:
+  • plugin call _julia, eval "x=1"
+The designated entry point to the plugin, the C++ function stata_call(), would then receive and process the two arguments after the comma, which tell it to have Julia evaluate x=1.
+Since calls to the plugin are in general arcane and sometimes require preparatory steps, the jl program serves as an essential and more user-friendly wrapper for it.
 That extra layer complicates access to local macros.
@@
 It will not appear in the context a step up, which could be your code.
-As a workaround, every time the Julia function st_local() is called, it adds the name of the macro it is writing to another local, __jllocals.
+As a workaround, every time the Julia function st_local() is called, it adds the name of the macro it is writing to another local in the name space of the jl program, called __jllocals.
 Before exiting, jl copies every macro in that list to its caller’s context using the undocumented Stata command c_local.
@@
 The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros.
-In partial compensation, single-line jl: commands can quote locals, because Stata processes the lines before passing them to jl.
-For example:
+As partial compensation, single-line jl: commands can quote locals, because Stata processes the lines before passing them to jl. For example:
   • local varname x1
@@
 This trick will not work in an interactive Julia session.
+[^14] The single, compound line jl: st_local("a","123"); print("‘a’") will also fail.
+The second command’s reference to a will be expanded before control is passed to Julia.
+The commands will work as intended if split into two lines.
 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
-The specification of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
+StataCorp’s definition of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
 One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double.
@@
 So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame.
-The jl commands can do nothing about the first conversion.
+The jl commands can do nothing about the first conversion since that is performed by Stata.
 But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision.
-This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive all the data, rather than a separate block for each of variously typed columns.
+This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
 Separately, Stata and Julia represent missing values in different ways.
@@
 Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values.
-It also allows observations of string variables to be missing.
-The julia package translates missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices).
+It thus distinguishes between empty and missing strings.
+The jl subcommands translate missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices).
 For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
-Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. The julia plugin contains optimized code that runs when such clauses are not specified.
+Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. When such clauses are not specified, the julia plugin runs specially optimized code.
 Data transfers are thus fastest when they map double precision to double precision and include no missing values or sample restrictions.
@@
 4.5 Simulated read-eval-print loop
-The Julia language system has many components—for interpreting and compiling code, managing memory, and so on.
+The Julia language system has many components—for compiling code, managing memory, and so on.
 The core parts are written in C++ and distributed as compiled libraries.
 Other components are themselves written in Julia, including the front end, which users often think of as Julia per se.
-The front end is usually called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat.
-But the REPL provides other conveniences: typing ”?” starts help mode; typing ”]” instead switches to the Julia package manager; some use is made of color; arrow keys allow one to explore one’s command history.
+The front end is called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat.
+But the REPL provides other conveniences.
+Typing ”?” starts help mode.
+Typing ”]” instead switches to the Julia package manager.
+Greek letters and other symbols can be entered through a TeX-like syntax.
+Arrow keys allow one to explore one’s command history.
 While there may be a way to run the full Julia REPL inside of Stata, the jl command does not do that.
@@
 A workhorse for jl is the Julia system’s jl_eval_string() function.
-It accepts a C string holding a Julia expression, evaluates it, and returns the result in a C structure that represents a generic Julia object.
+It accepts a C string holding a Julia expression, evaluates it, and returns the result in a C data structure that represents a generic Julia object.
 All Julia lines submitted to jl pass through this function.
-As its name suggests, the job of the function is to evaluate strings, not provide a user experience.
+As its name suggests, the function’s job is to evaluate strings, not provide a user experience.
 Especially when in interactive mode, jl takes steps to create an ersatz Julia REPL.
@@
 Another REPL feature that jl simulates matters for code such as this:
-  • jl
   • S = 0
@@
     print(S)
-    exit()
 This snippet prints the sum of the first 10 whole numbers—at least when run in the Julia REPL.
-Actually, for reasons having nothing to do with Stata, the snippet will crash if run non-interactively, as part of a stored program.
-Under Julia’s “soft scoping” rule, since S is not declared or initialized within the for loop before the first attempt to increment it, the attempt will trigger an undefined variable error.
+Actually, for reasons having nothing to do with Stata, the snippet will crash if run non-interactively, as part of a stored Julia program file.
+Under Julia’s “soft scoping” rule, since S is declared as a global variable outside the for loop or any other structure such as a function definition, it is invisible within the for loop.
+The first attempt to increment it from within the loop will trigger an undefined-variable error.
 Because that rule is counterintuitive, Julia’s REPL engineers an exception for interactive use.
 [^16] See the documentation at docs.julialang.org/en/v1/manual/variables-and-scoping/#local-scope.
-jl: imitates the exception by calling the Julia softscope() function on the user’s input after parsing it.
+jl imitates the exception by calling the Julia softscope() function on the user’s input after parsing it.
 softscope() is itself part of the code base of the Julia REPL.
+[^17] Thanks to Mark Kittisopikul for the idea of calling softscope().
 Still, jl’s REPL simulation is incomplete. Typing ? and ] alone have no effect. Plots do not automatically appear, which is why the examples in section 3.3 explicitly call display().
@@
 But it can add up in a program that calls jl: many times.
-The _jl: variant of the prefix command bypasses all of this processing.
+The _jl: variant of the prefix command bypasses all of this REPL simulation.
 5 The julia package
 5.1 Architecture
-The julia package consists of three intertwined pieces, all of which have been mentioned.
+The julia package consists of three intertwined pieces.
 The core is a shared library written in C++ according to the specification for Stata plugins.
 It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon.
-The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending commands to Julia and returning the results.
+The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
 The file “jl.ado” serves as a Stata front end for the plugin, providing a more intuitive interface and performing tasks better done in the ado language than in C++.
-At the suggestion of Jeffrey Pitblado at StataCorp, this program is called jl rather than julia because Stata might introduce an official julia command someday.
-When jl is first called in a Stata session, it briefly runs the Julia REPL through a shell command in order to query the location on disk of Julia’s shared libraries.
+At the suggestion of Jeffrey Pitblado at StataCorp, this program is called jl rather than julia because Stata might introduce an official julia command.
+When jl is first called in a Stata session, it briefly runs the true Julia REPL through a shell command in order to query the location on disk of Julia’s shared libraries.
 It passes the location to the plugin, which loads the libraries.
-This lets the plugin access Julia.
+[^18] Thanks to David Anthoff and Elliot Saba for guidance on this solution.
 jl also calls Stata’s findfile to track down the third key piece of the package.
-“stataplugininterface.jl” defines a module of Stata interface functions in Julia, such as st_local() and st_data().
-These work by calling entry points in the Stata plugin interface, which are exported to Julia by the plugin.
-(The Julia language includes a way to call C functions.) At start-up, jl therefore stores the location of “jl.plugin” in the stataplugininterface module.
+“stataplugininterface.jl” defines a module of Stata interface functions in Julia, such as st_numscalar() and st_matrix(), which are used in examples above.
+[^19] It would be more “Julianic” to name the file “StataPluginInterface.jl”.
+But rather than being posted to Julia’s general package registry, the file is distributed as part of a Stata package, and the Stata package installer cannot handle uppercase characters.
+These work by calling C-language entry points in the Stata plugin interface, which are exported to Julia by the plugin.
+(The Julia language includes a way to call C functions.) At start-up, jl therefore stores the location of “jl.plugin” in a variable in the loaded stataplugininterface module.
 5.2 Stata commands in the julia package
@@
 where juliaexpr is an expression to be evaluated in Julia.
-jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to set the upper limit on the number of threads:
+jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to cap the number of threads:
   • jl start [, threads(# | auto)]
 Specifying threads(auto) instructs Julia to use its own heuristic to pick the number of threads. Executing jl start is optional, for jl will initialize itself as needed.
-Several subcommands help manage Julia packages.
+Several subcommands help manage Julia packages, which are discrete libraries of code structured for easy sharing and are usually hosted on GitHub.
 Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments.
@@
     jl AddPkg name, [minver(X.Y.Z)]
-GetEnv displays the directory of the current package environment and returns it in r() macros. On my computer, which is running Julia 1.10, the command displays:
-  • . jl GetEnv
-[…]
-jl SetEnv moves to a chosen, named environment, which is merely a subdirectory of the default environment directory:
+GetEnv displays the directory and contents of the current package environment and returns them in r() macros.
+jl SetEnv moves to a chosen, named environment, which is merely a subdirectory of the default environment directory.
+If the environment does not exist, it is created and populated with the DataFrames.jl and CategoricalArrays.jl packages.
+One can return to the default environment with “SetEnv .” or just “SetEnv”.
+[^20] In distributed computing, Julia worker processes do not automatically inherit the master’s package environment.
+One can pass on the master’s environment through command line switches when launching the workers, such as in addprocs(4, exeflags="--project=\$(Base.active_project())").
+On my computer, creating a new environment called myenv produces this output:
   • . jl SetEnv myenv
 […]
-If the environment does not exist, it is created and populated with the DataFrames.jl and CategoricalArrays.jl packages.
-One can return to the default environment with SetEnv .
-or just SetEnv.
-[^15] In distributed computing, Julia worker processes do not automatically inherit the master’s package environment.
-One can pass on the master’s environment through command line switches when launching the workers, such as in addprocs(4, exeflags="--project=\$(Base.active_project())").
-Last in the package management family is the AddPkg option, which installs a package in the current environment or, if minver() is specified, updates it to the latest version if the current version falls below the minimum.
+[…]
+      [324d7699] CategoricalArrays v0.10.8
+      [a93c6f00] DataFrames v1.6.1
+The output ends with a list of currently installed packages—here, the two that are automatically installed.
+Last in the package management family is the AddPkg subcommand.
+Within the current environment, the subcommand installs a package; or, if minver() is specified, updates the package to the latest version if the current version falls below the specified minimum.
 Versions X.Y.Z are in the three-part semantic versioning format.
@@
 The first six listed below are meant more as programmer’s commands while the last two are higher-level.
+[^21] Johannes Boehm suggested adding high-level data-copying commands.
 jl GetVarsFromDF, for example, requires that the Stata data set already be at least as tall as the DataFrame from which it will receive variables.
 In contrast, jl use works like Stata’s use command: it replaces the current data set, and will fail unless the replace option is included, or the current data set has not changed since it was last saved.
-  • jl PutVarsToDF [varlist] [if] [in], [destination(string) cols(string) nolabel nomissing doubleonly] jl GetVarsFromDF varlist [if] [in], [cols(string) source(string) replace nomissing] jl PutVarsToMat [varlist] [if] [in], destination(string) jl GetVarsFromMat varlist [if] [in], source(string) jl PutMatToMat matname, [destination(string)] jl GetMatFromMat matname, [source(string)] jl save [dataframename], [nolabel nomissing doubleonly] jl use dataframename, [clear] jl use varlist using dataframename, [clear]
+  • jl PutVarsToDF [varlist] [if] [in], [destination(string) cols(string) nolabel nomissing doubleonly] jl GetVarsFromDF varlist [if] [in], [cols(string) source(string) replace nomissing] jl PutVarsToMat [varlist] [if] [in], destination(string) jl GetVarsFromMat varlist [if] [in], source(string) [replace] jl PutMatToMat matname, [destination(string)] jl GetMatFromMat matname, [source(string)] jl save [dataframename], [nolabel nomissing doubleonly] jl use dataframename, [clear] jl use varlist using dataframename, [clear]
 The varlists and matnames before the commas refer to Stata variables or matrices.
 If an optional varlist is omitted, it defaults to *, meaning all variables in the current data set in their current order.
-A destination() or source() option after the comma refers to a Julia matrix or DataFrame.
+A destination() or source() option after the comma should contain the name of a Julia matrix or DataFrame.
 When an optional DataFrame name is not provided, it defaults to df.
@@
 Through the view, Julia code can read and write Stata variables while treating them as part of a matrix.
-The full list of functions:
+The full list of Stata-interfacing Julia functions:
 SF_nobs() Number of observations in Stata data set
@@
 SF_var_is_strl(i) Whether variable i is a strL
-SF_var_is_binary(i, j) Whether observation i of variable j is a binary strL
+SF_var_is_binary(j, i) Whether observation j of variable i is a binary strL
 SF_sdatalen(j, i) String length of variable i, observation j
 SF_is_missing() Whether a Float64 value is Stata missing
-SV_missval() Stata floating-point value for missing
+SF_missval() Stata floating-point value for missing
 SF_vstore(j, i, val) Set observation j of variable i to val (numeric)
@@
 SF_macro_save(mac, tosave) Set macro mac
-SF_macro_use(mac, maxlen) First maxlen characters of macro mac
+SF_macro_use(mac) Contents of macro mac
 SF_scal_save(scal, val) Set value of scalar scal
@@
 st_matrix(matname) Get numeric Stata matrix
-st_matrix(matname, jlmat) Put Julia matrix in pre-existing Stata matrix
+st_matrix(matname, jlmat) Put Julia matrix in existing Stata matrix
 st_data(varnames) Get Stata variables, as matrix
@@
       sample = rand(Bool, st_nobs())    # random Boolean vector defining subsample
-      v = st_view("price mpg", sample)  # view onto subsample of two vars
-      v ./= 2                           # halve these Stata data points
+      v = st_view("price mpg", sample)  # view onto subsample of two Stata variables
+      v ./= 2                           # halve these values in the Stata data set
       st_numscalar("s", sum(v))         # sum the halved data into Stata scalar s
@@
 6 Conclusions
-Julia is a strong option for universal, back-end development of statistical applications.
-Working in Julia comes with some drawbacks, including the challenge of learning a complex new language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation.
+Julia is a strong option for universal back-end development of statistical applications.
+Working in Julia comes with some drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation.
 But Julia and nearly all Julia packages are free and open and run on all popular operating systems.
@@
 Instead of implementing new methods separately in Stata or R or Python, or in a low-level language such as C, developers can write the back ends once, in Julia.
-Here, WildBootTests.jl shows the way: it is an optional back end for boottest in Stata and a required back end for instrumental variables inference with fwildclusterboot in R.
+Here, WildBootTests.jl shows the way: it is an optional back end for boottest in Stata and fwildclusterboot in R, and a required back end for instrumental variables inference with the latter.
 For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia.
@@
 Most features needed to make Stata objects accessible in Julia can and have been implemented.
-The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed, new features, or more flexible syntax (as for plotting).
-reghdfejl is a perhaps the “killer app” for the Julia-Stata link.
-It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for reghdfe.
-And reghdfejl preliminarily handles nonlinear HDFE models, which brings new functionality to Stata users.
+The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed or new features.
+reghdfejl might be the “killer app” for the Julia-Stata link.
+It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the older program.
+It also handily outpaces the newly augmented areg.
+And, preliminarily, reghdfejl handles nonlinear HDFE models, thus bringing new functionality to Stata users.
 StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java.
@@
 They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols.
-They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it first.
-They could make Mata interoperate with Julia.
+They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it one observation at a time, or even without copying it at all.
+They could give Mata the ability to call Julia functions.
 The goal for the julia package was to prove a concept. In that I think it has succeeded.
+7 Acknowledgments
+I thank Sergio Correia, Páll Haraldsson, and an anonymous reviewer for comments.

### v4

@@
 But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms.
-For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can both call the same Julia back end.
-And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls a Julia package for tenfold acceleration on hard problems.
+For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can use the same Julia back end.
+And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects while calling a Julia package for tenfold acceleration on hard problems.
 reghdfejl also supports nonlinear fixed-effect models that cannot otherwise be fit in Stata—though preliminarily, as the Julia package for that purpose is immature.
@@
 1 Introduction
-Whenever you issue a command to perform a statistical calculation, you initiate a chain of events of bewildering complexity: your request cascades through layers of code crafted over decades by thousands of programmers in multiple languages, as it is expanded into millions or billions of computational steps—steps that are executed, if you’re lucky, in the time it takes to read this sentence.
+Whenever you issue a command to perform a statistical calculation, you initiate a chain of events of unfathomable complexity: your request cascades through layers of code crafted over decades by thousands of programmers in multiple languages, as it is expanded into millions or billions of computational steps—steps that are executed, if you’re lucky, in the time it takes to read this sentence.
 In Stata, for example, a call to regress or gsem starts out in the ado programming language.
@@
 In modern Stata, an ado program can even call Python or Java.
-The called code will probably invoke the LAPACK library, which is written in Fortran, to invert matrices or solve linear systems.
+The called code will probably invoke the Fortran-based LAPACK library to invert matrices or solve linear systems.
 Here I introduce a package that forges a fresh link between Stata and another language, Julia.
@@
 More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it.
-Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution.
-(Python and C++ is a common two-language combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse.
+Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development for speed of execution.
+Python and C++ is a common two-language combination, as is R and C++.
+Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse.
 A final strength is that, like Lisp, Julia can treat code as data.
-This feature opens the door to meta-programs that can automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
+This feature opens the door to meta-programs that can automatically optimize one’s code by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
 Julia has disadvantages.
-While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented.
+While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit common models, and generate plots—many packages are immature or poorly documented.
 Standards for storing and returning results are less well developed than in Stata.
 Much less statistical functionality is available than in Stata or R, or even Python.
-Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development.
+Since Julia is a great home for numerical software, but not (yet) a great home for the end user, a sweet spot for the language is back end development.
 Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms.
-The core work can augmented and optimized in one place, rather than separately on each platform.
+The core can be augmented and optimized in one place, rather than separately on each platform.
 The julia package links Stata to Julia.
@@
 StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003.
-It added Mata in Stata 9.
+It added the built-in Mata language in Stata 9.
 Fiedler (2012) envisioned running Python inside of Stata.
@@
 Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk.
-Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (JuliaConnectoR) packages link R to Julia.
+Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (Lenz et al., 2022) packages link R to Julia.
 Another package named JuliaCall does the same for Python.
@@
 From the user’s point of view, the package has three components:
-  • The jl: command for Stata, which prefixes single Julia lines or, when typed by itself (with or without the colon), starts a Julia session in Stata.
+  • The jl command for Stata, which prefixes single Julia lines or, when typed by itself, starts a Julia session in Stata.
   • jl subcommands for managing Julia packages and transferring data between Stata and Julia.
   • Julia functions to read and write Stata macros, scalars, matrices, and variables.
-The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package for Stata.
+The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started with Julia requires nothing more than installing the julia package for Stata.
 Two user-written Stata packages call Julia via julia.
-The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata implementation I originally wrote for boottest.
-And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016).
+The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata back end I originally wrote for boottest.
+And the new reghdfejl mimics the reghdfe and ivreghdfe programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016).
 reghdfejl calls the Julia package FixedEffectModels.jl, whose lead developer is Matthieu Gomez.
@@
 It starts by demonstrating how to run Julia from Stata, in section 2.
-Section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia.
+Section 3 introduces applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia.
 Section 4 gives tips and warns of some limitations.
@@
 2.1 Installation and first examples
-Julia is an interpreted language in the sense that you can only run programs written in it if you have installed it.
-[^2] PackageCompiler.jl can build stand-alone Julia apps, but the files take hundreds of megabytes because they contain complete copies of Julia and its libraries.
-In contrast, compiled C++ and Fortran programs run on computers without a C++ or Fortran compiler.
-The preferred way to get Julia is to download the juliaup installation manager.
-You do that in Windows by installing Julia (not juliaup) from the Windows Store.
-[^3] The instructions might change.
-See github.com/JuliaLang/juliaup#installation for the latest.
-In Linux and macOS, you can instead start Stata and type
-  • ! curl -fsSL https://install.julialang.org | sh -s -- -y
-(including the “!” at the beginning). In Linux, you should then restart Stata and the shell to trigger the addition of Julia’s location to the PATH environment variable.
 To install the julia package in Stata, type:
@@
   • . jl: "Hello world!"
+      (output omitted )
     Starting Julia
@@
 This operation may take several minutes.
-In Windows, many mysterious console windows may appear and disappear as required libraries are downloaded and pre-compiled.
-In macOS and Linux, output from the installation process may appear in terminal windows.
-For on first use, jl installs two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on.
-DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, i.e., rectangular grids whose columns have names and potentially different data types.
-CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
+In Windows, mysterious console windows may appear and disappear.
+Julia and the Julia version manager will be installed, if they haven’t been already.
+The version manager will be configured to guarantee that Julia 1.11, the version currently required by jl, is always available.
+jl will also install two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on.
+DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, meaning rectangular grids whose columns have names and potentially different data types.
+CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” (Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.)
+While Julia’s developers work hard to assure that programs written in older versions work in newer ones too, they do not guarantee backward compatibility in Julia’s low-level, C-based interface, which is what the julia package for Stata links to.
+Changes to that interface can cause jl to completely crash Stata.
+(Ask me how I know.) This is why jl is designed to work only with the latest version of Julia at this writing, 1.11.
+Fortunately, Julia’s version manager makes it easy to maintain several Julia versions at once on a computer.
+So if you want to use a different version outside of Stata, jl’s need for 1.11 won’t interfere.
+[^2] If revisions to Julia’s innards break the julia package for Stata, I expect to revise the package to stay up-to-date.
+The required Julia version may change.
 In general when using Julia, you will experience delays at four points: when packages are installed; when they are first used; when they are first used in a Julia session; and when particular features within them are first used, such as clustering standard errors.
 A thrust of recent development in the Julia language has been to move the behind-the-scenes processing to earlier stages in order to shorten the lags during regular use.
-In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them, then copy the data to a Julia DataFrame named auto:
+In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them in Stata, make sure the package I will use for estimation in Julia is installed on my computer, and then copy the data to a Julia DataFrame named auto:
   • . sysuse auto
@@
             _cons     13743.64     2176.8     6.31   0.000      9150.99    18336.28
+    . jl AddPkg FixedEffectModels
     . jl save auto
     Data saved to DataFrame auto in Julia
-jl save is not a Julia command.
-It is a subcommand of the Stata command jl, which is part of the julia package.
+jl AddPkg and jl save are not Julia commands.
+They are subcommands of the Stata command jl, which is part of the julia package.
 If an object named auto already exists in the Julia environment, it is overwritten by the incoming copy of the Stata variables.
@@
 The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing.
-By default, jl save made all the columns accept missing.
-jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector.
-While this Julia version of the variable is also stored as a column of integers—1’s and 2’s—it presents as a text variable.
-[^4] Julia categorical variables are not quite comparable to Stata variables with data labels.
-Commands involving the Julia variable refer to display values such as “Domestic,” not to numeric codes.
+By default, jl save makes all the columns accept missing.
+jl save also converts the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector.
 The type name “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
+While the Julia version of foreign is also represented internally as a column of integers, it presents as a text variable.
+All commands involving the Julia variable would refer to display values such as “Domestic,” not to numeric codes as in Stata.
 Next I run the same regression as before, but now in Julia, with the reg() function of the Julia package FixedEffectModels.jl:
-  • jl> . jl AddPkg FixedEffectModels  # assure FixedffectModels installed
-    jl> . using FixedEffectModels      # load it for use
+  • jl> . using FixedEffectModels      # load package for use
     jl> . reg(auto, term(:price) ~ term(:headroom) + term(:foreign)&term(:mpg),
@@
     (Intercept)              13743.6      2176.8     6.31369     <1e-04   9129.03   18358.2
-The Stata and Julia results match. But that Julia regression command needs explaining. First some cosmetic comments:
-  • Because the regression command was long, I broke it in two.
-After I hit the enter key at the end of the first line, jl detected that the line was not grammatical by itself and printed “...” to prompt me to continue.
+The Stata and Julia results match. But the Julia regression command needs explaining. First some cosmetic comments:
+  • Because the regression command is long, I break it in two.
+After I hit the enter key at the end of the first line, jl detects that the line is not grammatical by itself and prints “...” to prompt me to continue.
   • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears above.
@@
 In the example, an entire FormulaTerm is passed as the second argument to reg().
-In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
+In this way, the primitives of a flexible programming language are composed to articulate a statistical model.
 This standard for representing regression models is not intrinsic to Julia.
@@
 StatsModels.jl defines a macro called @formula, which lets one write models more simply.
-Thus:
+Thus, I can write:
   • jl> . m = reg(auto, @formula(price ~ headroom + foreign&mpg), Vcov.cluster(:turn))
-    [identical output omitted]
+      (output omitted )
+This is how regressions are usually specified in Julia.
 Before this new command is compiled and executed, @formula turns it into something closer to the earlier regression command.
@@
 We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()).
-But the interactive mode is not accessible to do files.
+But the interactive mode is not accessible in do files.
 To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line.
 It is typographically, not syntactically, split across multiple lines with Stata’s continuation token, ///.
-When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
+When putting many Julia commands on what is syntactically one line, most Julia commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
 The code for this new Julia function, XQX(), is mostly straightforward.
@@
 Allocating and deallocating memory itself takes time.
-Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory.
+Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are transferred to and from the computer’s main memory.
 In contrast, the triply-looped Julia code demands essentially no temporary storage.
 This inefficiency of Mata and Python can be seen as an artifact of the two-language problem Julia was designed to solve.
-The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are nearly impossible to revise, yet not optimized for the purpose at hand.
+The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are not optimized for the purpose at hand and nearly impossible to revise.
 To create a test bed for these functions, I construct ${\bf X}$ as a Stata data set and ${\bf Q}$ as a Stata matrix, then copy them into Mata, Python, and Julia.
@@
 The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because that is when it is compiled down to machine code.
-On a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python.
+We see that on a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python.
 In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
 To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above.
-The Julia package manager would run the script after installing the package in order to trigger immediate compilation.
+The Julia package manager would automatically run the script after installing the package in order to trigger immediate compilation.
 Julia would save the generated machine code for future use.
@@
 While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007).
-As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic.
+As a result, it can normally only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic.
 reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and then transfer control to ivreg2.
@@
   • vce(bs, reps(1000) seed(42) procs(4) cluster(year))
-specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from groups defined by year.
+specifies that 1000 bootstrap replications be performed, split across four CPU threads running in parallel, with the bootstrap data drawn groupwise from groups defined by year.
 One could attain the same results with bootstrap:reghdfe or bs:reghdfe.
 But reghdfejl implements vce(bs) in a radically faster way.
-After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times in a bootstrap sample, and so on.).
-To demonstrate reghdfejl’s speed, I create a data set with 10 million rows and two sets of fixed effects.
+It copies the primary data just once to Julia.
+It effects the resampling merely by updating a weight variable, so that, for example, a cluster’s weights are doubled if it is drawn two times in a bootstrap sample.
+To demonstrate reghdfejl’s speed on hard problems, I create a data set with 10 million rows and two sets of fixed effects.
 I run the same regression with reghdfe and reghdfejl, as well as with the built-in areg command, which in StataNow 18.5 gained the ability to absorb multiple fixed effects.
-While reghdfe and reghdfejl can cluster standard errors on multiple variables (CGM_2011), areg cannot, so the example is restricted to one-way clustering:
+While reghdfe and reghdfejl can cluster standard errors on multiple variables (Cameron et al., 2011), areg cannot, so the example is restricted to one-way clustering:
   • set obs 10000000
@@
 reghdfejl takes just 1.9 seconds.
-But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 12 worker copies of Julia and triggers some just-in-time compilation in each.
-The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers reqested by the procs() suboption of the vce() option does not rise further.
 The reghdfejl package includes a partialhdfejl command for absorbing (partialling) fixed effects out of other variables.
@@
 Reducing the data to these products early in the calculation complicates the algebra the code must implement, but speeds execution when clusters are few.
-And the functionality in boottest, so revised, has been transplanted to the Julia package WildBootTests.jl.
+And the functionality in boottest, so revised, has been transplanted to Julia, in the package WildBootTests.jl.
 Adding the julia option to a boottest command line switches to the new back end.
@@
 [^8] Data and preparation code are available at davidroodman.com/david/Levitt.zip.
-Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for calling boottest to perform inference about changes in the size of the prisoner population affect crime rates:
+Here is a version of that example, which substitutes reghdfejl for ivregress and sets the stage for calling boottest.
+The purpose remains to perform inference about how changes in the size of the prisoner population affect crime rates:
   • use Levitt, clear
@@
         gridmin(-2) gridmax(2) nograph
-On a single CPU core, version 2.3.5 of boottest, the one published with Roodman et al.
-(2019), needs 210 seconds to perform the two boottest calls.
+On a single CPU core the version of boottest published with Roodman et al.
+(2019) (version 2.3.5), needs 210 seconds to perform the two boottest calls.
 The latest version finishes in 7.4 seconds, even without calling Julia.
 Appending julia to the boottest line above lowers the time to 4.1 seconds.
-(However, the first time boottest is called after installation, there is a significant delay for pre-compilation, just as when jl is first run, as described in section section 2.1.)
 The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE, and optionally does so to perform the wild bootstrap after ordinary least squares.
@@
 To indicate their potential for Stata users, I generate two graphs with the plotting package Makie.jl.
-Makie works with several plotting back ends, which render output in formats designed for different applications, such as interactive graphs in web pages and publication-quality image files.
+Makie works with several plotting back ends, which render output in formats suited to different applications, such as interactive graphs in web pages and publication-quality images.
 The Cairo back end is best for the latter.
 It is bundled with Makie in CairoMakie.jl:
-  • jl AddPkg CairoMakie  # one-time installation
-    jl: using CairoMakie  # use now
-In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.111111blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command.
+  • jl AddPkg CairoMakie  // one-time installation
+    jl: using CairoMakie  // use now
+In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.999blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command.
 margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights.
@@
       f |> display  # equivalent to display(f)
-      save("fig1.png", f)
+      save("fig1.pdf", f)
     exit()
-This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines.
+This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window.
+That is, I drop the “;” and “///” from the ends of lines.
 The first Julia line imports the data as a matrix, converts it to a DataFrame, and names the columns.
@@
 That line takes advantage of Julia’s piping syntax for function calls, using the symbol “|>”.
-(display(f) would be equivalent.) Last, the image is saved to disk.
+(display(f) is equivalent.) Last, the image is saved to disk.
 It appears in Figure 1.
@@
 One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other.
-reghdfejl’s vce(bs) option uses that feature.
 A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space.
-That is the sort @tturbo exploits.
-The maximum number of threads Julia allows is set when it launches, and it can default to 1.
+That is the sort @tturbo and the vce(bs) option of reghdfejl exploit.
+With regard to the latter, the maximum number of threads allowed in each Julia instance is set when the instance launches.
+It can default to 1.
+In Stata, you can determine the current limit with the command jl:Threads.nthreads().
 A good guess for the optimum is the number of “performance” or “big” cores on the CPU.
-You can determine the current limit with the command jl:Threads.nthreads().
-How best to set Julia’s thread limit depends on the operating system.
+How to set Julia’s thread limit depends on the operating system.
 In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS.
@@
 The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment.
-For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from a Julia matrix.
+For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from Julia.
 A subtler limitation pertains to local macros.
@@
 Before exiting, jl copies every macro in that list to its caller’s context using the undocumented Stata command c_local.
-However, the same complication prevents one from reading Stata locals from Julia.
+However, the same complication prevents one from reading your Stata locals from Julia.
 The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros.
@@
     jl: sum(df.`varname')
-This trick will not work in an interactive Julia session.
+This trick does not work in interactive Julia sessions (ones started by typing jl by itself).
 [^12] The single, compound line jl: st_local("a","123"); print("‘a’") will also fail.
@@
 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
-StataCorp’s definition of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
-One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double.
+Certain details of the Stata plugin interface introduce wrinkles into the copying of data between Stata and Julia.
+One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only truck in double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double.
 By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types.
@@
 But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision.
-This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
+This typically saves time, by reducing conversions and allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
 Separately, Stata and Julia represent missing values in different ways.
@@
 The jl subcommands translate missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices).
-For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
-Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. When such clauses are not specified, the julia plugin runs specially optimized code.
+For data known to contain no missing values, adding the nomissing option to the copying commands will save time by skipping this step.
+Finally, the SPI lets plugin calls include if or in clauses to restrict the Stata sample on which the commands will operate.
+When such clauses are not specified, the julia plugin runs specially optimized code.
 Data transfers are thus fastest when they map double precision to double precision and include no missing values or sample restrictions.
-Consider this example, which creates a 10 million $\times$ 10 set of double-precision numbers and copies them to a Julia DataFrame named demo:
+Consider this example, which creates a 100 million $\times$ 10 set of double-precision numbers and copies them to a Julia DataFrame named demo:
   • set obs 100000000
@@
 Other components are themselves written in Julia, including the front end, which users often think of as Julia per se.
-The front end is called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat.
+The front end is called the REPL, for “Read-Eval-Print Loop.” As its name makes plain, the REPL’s main job is to read in commands, evaluate them, print the results, and repeat.
 But the REPL provides other conveniences.
@@
 But it can add up in a program that calls jl: many times.
-The _jl: variant of the prefix command bypasses all of this REPL simulation.
+The _jl: variant of the prefix command saves time by bypassing all of this REPL simulation.
 5 The julia package
@@
 It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon.
-The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
+At installation, the “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
 The file “jl.ado” serves as a Stata front end for the plugin, providing a more intuitive interface and performing tasks better done in the ado language than in C++.
@@
 where juliaexpr is an expression to be evaluated in Julia.
-jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to cap the number of threads:
+jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to set the number of available threads:
   • jl start [, threads(# | auto)]
@@
 Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments.
-Julia also makes it easy to change the current environment.
+Julia also lets you create new environments and switch among them.
 This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration.
@@
   • . jl SetEnv myenv
-[…]
-[…]
+    Current environment: myenv, at /Users/davidroodman/.julia/environments/v1.11/myenv
+    Status ´~/.julia/environments/v1.11/myenv/Project.toml´
     [324d7699] CategoricalArrays v0.10.8
-      [a93c6f00] DataFrames v1.6.1
-The output ends with a list of currently installed packages—here, the two that are automatically installed.
+    [a93c6f00] DataFrames v1.7.0
 Last in the package management family is the AddPkg subcommand.
@@
 5.3 Julia functions in the julia package
-Functions in the stataplugininterface module can also be characterized as low- or high-level.
+Functions in the stataplugininterface.jl module can also be characterized as low- or high-level.
 The low-level ones wrap and mimic functions in the Stata plugin interface.
-For example, SF_vdata() extracts a single observation on a single variable.
+For example, SF_vdata() extracts a single observation on a single Stata variable.
 The higher-level ones build on the low-level ones in order to imitate some of Mata’s Stata interface functions.
@@
 Through the view, Julia code can read and write Stata variables while treating them as part of a matrix.
-The full list of Stata-interfacing Julia functions:
+The full list of Stata-interfacing Julia functions follows:
 SF_nobs() Number of observations in Stata data set
@@
 Julia is a strong option for universal back-end development of statistical applications.
-Working in Julia comes with some drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation.
-But Julia and nearly all Julia packages are free and open and run on all popular operating systems.
+True, working in Julia comes with drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation.
+But Julia and nearly all Julia packages are free and open and run on all popular platforms.
 The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain.
@@
 reghdfejl might be the “killer app” for the Julia-Stata link.
-It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the older program.
-It also handily outpaces the newly augmented areg.
+It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the latter.
+It also outpaces the newly augmented areg.
 And, preliminarily, reghdfejl handles nonlinear HDFE models, thus bringing new functionality to Stata users.
-StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java.
+StataCorp could tighten Stata’s integration with Julia, as it has done with Python and Java.
 Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables.